### v2

# Is Large-scale Pretraining the Secret to Good Domain Generalization?

Abstract
Multi-Source Domain Generalization (DG) is the task of training on multiple source domains and achieving high classification performance on unseen target domains. Recent methods combine robust features from web-scale pretrained backbones with new features learned from source data, and this has dramatically improved benchmark results. However, it remains unclear if DG finetuning methods are becoming better over time, or if improved benchmark performance is simply an artifact of stronger pre-training. Prior studies have shown that perceptual similarity to pre-training data correlates with zero-shot performance, but we find the effect limited in the DG setting. Instead, we posit that having perceptually similar data in pretraining is not enough; and that it is how well these data were learned that determines performance. This leads us to introduce the Alignment Hypothesis, which states that the final DG performance will be high if and only if alignment of image and class label text embeddings is high. Our experiments confirm the Alignment Hypothesis is true, and we use it as an analysis tool of existing DG methods evaluated on DomainBed datasets by splitting evaluation data into In-pretraining (IP) and Out-of-pretraining (OOP). We show that all evaluated DG methods struggle on DomainBed-OOP, while recent methods excel on DomainBed-IP. Put together, our findings highlight the need for DG methods which can generalize beyond pretraining alignment.
1 Introduction
Domain Generalization (DG) addresses the challenge of enabling AI models to generalize from known domains to unseen ones, a critical task given the inevitable distribution shifts between training and real-world deployment (Saenko et al., 2010). DG pipelines typically consist of three stages: pretraining a model on a large, general dataset; finetuning the model with one or more source domains; and finally evaluating the model on target domains that are distinct from source domains. More and more, DG methods rely on huge-scale foundation models for initialization. Simultaneously, finetuning has increasingly incorporated regularization to prevent catastrophic forgetting. As a result, it remains unclear whether DG adaptation methods are genuinely improving or if enhanced benchmark performance is simply due to stronger pre-training combined with regularization, or even the presence of target domains within the hundred million-scale pre-training data.
In this work, we examine the reliance of recent DG methods on pre-trained features. We focus on CLIP-based models, as they are used in all state-of-the-art DG methods (Addepalli et al., 2024; Cho et al., 2023; Shu et al., 2023; Mao et al., 2024). While prior studies  (Mayilvahanan et al., 2024) have shown that perceptual similarity to pre-training data explains zero-shot performance—referred to as the Image Similarity Hypothesis—we find this relationship to be limited in the DG setting. Despite evidence of target domains being present in pre-training (Figure 6), perceptual similarity alone does not fully explain accuracy in the DG context. We propose that it is not just the presence of similar data in pre-training that matters, but also how well this data was learned. To this end, we introduce the Alignment Hypothesis, which states that pre-trained alignment between image and class embeddings is still predictive of DG performance even after source finetuning. We find that, performance for low alignment samples can be almost 0, while performance for high alignment samples is close to perfect. These results confirm the Alignment Hypothesis. As illustrated in Figure 1, these findings suggest that current DG methods largely fail to learn new, general features from the source data when the pretraining does not already provide a strong alignment.
Figure 1: An overview of desired and actual behaviour of DG methods. 1) DG methods are initialized with foundation models like CLIP. Pre-trained embeddings are relatively well aligned with ground truth labels on both source and target data for most samples (In-Pretraining, IP), but some samples are not well aligned (Out-of-pretraining, OOP). 2) An ideal DG method would strengthen alignment for both OOP and IP data with ground truth labels. 3) Our analysis shows that DG methods only result in strong alignment for IP data, leaving OOP data misaligned (Figure 2).
The confirmation of the Alignment Hypothesis gives us a tool to separate in-pretraining (IP) data from out-of-pretraining (OOP) data for a particular backbone, and we do so for five DG datasets with OpenCLIP-ViT/B-16. We call the resulting splits DomainBed-IP and DomainBed-OOP. Evaluating on DomainBed-IP/OOP offers a view of where current DG methods fail and where they succeed. We find that all methods, including those considered state-of-the-art, perform poorly on OOP data, i.e. data that the pretrained backbone hadn’t already aligned well. Furthermore, recent state-of-the-art methods do not outperform older methods on OOP data. For example, CLIPood  (Shu et al., 2023) slightly under performs a combination of older methods (MIRO (Cha et al., 2022) + MPA (Arpit et al., 2022)) on DomainBed-OOP. At the same time, existing DG methods show exceptional performance on DomainBed-IP, sometimes even outperforming an oracle model trained on the target domain. These results suggest that future research should aim to enhance DG methods on low-alignment data while preserving the already strong performance on high-alignment data. In summary, we make the following contributions:
  • Introduce the Alignment Hypothesis: We demonstrate that pre-training alignment between image and class text embeddings is a stronger predictor of Domain Generalization (DG) success than the previously proposed Image Similarity Hypothesis  (Mayilvahanan et al., 2024). Based on this, we define In-Pretraining (IP) as data well-aligned with pre-trained embeddings, and Out-of-Pretraining (OOP) as data with weaker alignment.
  • Propose a new IP/OOP evaluation framework: We demonstrate that splitting target data by its alignment with the pre-trained backbone can effectively test Out-of-Pretraining (OOP) generalization. We will release IP/OOP splits for the DomainBed dataset to support future research.
  • Expose strengths and limitations of state-of-the-art DG methods: Using DomainBed-IP/OOP we find that leading DG methods perform well on data well-aligned by pre-training but struggle on misaligned samples, emphasizing the need for methods that move beyond reliance on pre-training.
2 Related Work
Multi-Source Domain Generalization: Domain Generalization aims to mitigate the impacts of domain shifts between source (training) and target (deployment) domains. One standard approach is domain-invariant feature learning, which leverages domain labels to learn domain-invariant features. CORAL  (Sun & Saenko, 2016) aligns second-order statistics, while DANN  (Ganin et al., 2016) uses an adversarial loss.  Gulrajani & Lopez-Paz (2020) show that ERM, which does not align features between domains, can outperform most prior work while being easier to tune. Another common approach is domain-aware data augmentation to expand the training domain to become closer to or even overlap the target domain. Inter-domain mixup  (Yan et al., 2020) blends images from different domains. Similarly, style transfer can diversify training images  (Zhong et al., 2022). Deep ensembles are effective for domain generalization (Arpit et al., 2022). Since they are computationally inefficient for inference, many recent works average model weights from either multiple finetuning runs or from a single training trajectory (Cha et al., 2021; Arpit et al., 2022; Rame et al., 2022; Jain et al., 2023; Li et al., 2023; Shu et al., 2023). More recently, several methods perform regularized finetuning towards the initialization of a pretrained model. This works under the assumption that pretrained features are useful for target data, and should not be unlearned. The general idea can be applied to weight space (L2SP (Xuhong et al., 2018)), feature space (MIRO (Cha et al., 2022)), or output space (CAR-FT (Mao et al., 2024),CLIPood (Shu et al., 2023)).
Large-Scale Pretraining for DG: Recent DG literature (Cho et al., 2023; Cha et al., 2022; Addepalli et al., 2024; Mao et al., 2024; Arpit et al., 2022) leverages large-scale pretrained initializations stronger than ImageNet (Russakovsky et al., 2015), and CLIP (Radford et al., 2021) is the most common choice. CLIP leverages a cross-domain contrastive loss to align images and captions. Due to the large scale of training data (typically at least 400 million samples) and the free-form nature of the text, CLIP enables effective zero-shot classification and learns features that generalize very well. Other choices for very strong pretraining include SWAG (Singh et al., 2022) and DinoV2 (Oquab et al., 2023). SWAG uses supervision from Instagram hashtags, while DinoV2 is trained without text supervision and instead relies on augmentation-based alignment. While our analysis focuses on image-text models like CLIP due to its popularity, the concept of alignment can extend to other types of pretraining models. We leave the exploration of this extension to future work.
Impact of Data on Model Performance: Several recent studies have explored the influence of pretraining data on model performance. Mayilvahanan et al. (2024) investigated how the presence of perceptually similar images in CLIP (Radford et al., 2021) pretraining affects performance, introducing the Similarity Hypothesis, which posits that nearest neighbor similarity is strongly correlated with zero-shot accuracy. Udandarao et al. (2024) demonstrated that concept frequency in pretraining is correlated with zero-shot performance and introduced a dataset focusing on infrequent concepts. Fang et al. (2022) found that diversity in pretraining data is critical for improving performance on benchmarks such as ImageNetV2 (Recht et al., 2019), ImageNet-R (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), and ObjectNet (Barbu et al., 2019). However, these studies focus on the zero-shot setting, where models are evaluated without further training. In contrast, we examine the domain generalization setting, where pre-trained models are fine-tuned on source domains and tested on held-out target domains. Our findings suggest that comparing target images to pre-trained images, as proposed by Mayilvahanan et al. (2024), is less predictive of final DG performance than directly measuring the alignment between the image and its class embedding.
3 Analyzing the role of pretraining in Domain Generalization
This work explores Multi-Source Domain Generalization for classification, where samples from multiple source domains (e.g., sketches, product photos) and a held-out target domain (e.g., wildlife camera images) are annotated with both domain and class labels. We construct a training dataset by aggregating all sample-label pairs from all training domains $d\in\{d_{1},\ldots,d_{n}\}$, denoted as
$D=\{(X^{d_{1}},Y^{d_{1}}),\ldots,(X^{d_{n}},Y^{d_{n}})\}.$
We initialize a classifier $f$ with a contrastively pre-trained vision-language model (e.g., CLIP) and finetune it on $D$. The scale of pre-training datasets is many orders of magnitude larger than that of source datasets. Most methods fully fine-tune $f$, though LP-FT (Kumar et al., 2021) fine-tunes the linear probe before the main network and Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022) only tunes attention layers. The performance is then evaluated on a held-out testing domain $d_{\text{test}}$. The key assumption is that $d_{test}$ has a different distribution from the source domains.
We aim to analyze how reliant existing DG methods are on pre-training. A recent analysis of CLIP proposed the Image Similarity Hypothesis  (Mayilvahanan et al., 2024), which supposes that high CLIP performance on a given test sample is a result of highly similar nearest-neighbor images in pre-training, and tested it on zero-shot classification tasks. They found a strong correlation between nearest-neighbor similarity and zero-shot classification performance, but did not analyze OOD performance after fine-tuning. Therefore, we apply an equivalent testing setup for the DG setting, where a pre-trained model is fine-tuned on a source distribution and tested on a different target distribution. We find only a limited influence of image similarity in Section 3.1. To better understand the role of pretraining in domain generalization, we introduce the Alignment Hypothesis, which we explore in detail in Section 3.2. We later use the Alignment Hypothesis to split DG datasets and analyze existing DG methods (Section 4).
3.1 Testing the Image Similarity Hypothesis
The Image Similarity Hypothesis  (Mayilvahanan et al., 2024) posits that test performance improves when there are perceptually similar images in the pre-training dataset. The PerceptualSimilarityScore measures perceptual similarity and is defined as the cosine similarity between a target image $I$ and its nearest neighbor $I_{k}$ in pre-training:
[…]
Figure 2: Comparing the Predictive Power of the Alignment and Image Similarity Hypotheses for Domain Generalization (DG). a.) Alignment Hypothesis: The cosine similarity between image and ground truth text-label embeddings before pre-training (Alignment Score) is highly predictive of model accuracy after fine-tuning on five DG datasets, with Alignment Score distributions shown in the colored histograms. This suggests that image-text pairs well-aligned during pre-training result in better performance on target tasks. b.) Image Similarity Hypothesis: In contrast, the cosine similarity between a test image and its closest match from the pre-training set (Perceptual Similarity Score) shows weaker predictive power for accuracy, implying that visual resemblance alone is less indicative of downstream performance.
[…]
3.2 Introducing the Alignment Hypothesis
To find a stronger predictor of DG accuracy than perceptual similarity, we focus on how effectively pre-training captures the relationship between an image and its label. This leads us to propose the Alignment Hypothesis, which states that if an input image and its corresponding text label (e.g., ‘A photo of a {cls}’) are well-aligned in the embedding space, final DG performance will be high. Crucially, alignment is measured before source fine-tuning while DG performance is measured after adaptation. This allows us to isolate the contribution of fine-tuning. Since models like CLIP optimize image-text pairs using a contrastive loss, cosine similarity between image and text embeddings is an alignment measure well coupled to their training objective. Therefore, we use it as our metric of pre-training generalization. More formally:
$\text{AlignmentScore}(I,T)=\frac{\langle f_{I}(I),f_{T}(T)\rangle}{\|f_{I}(I)\|\cdot\|f_{T}(T)\|}$
where $f_{I}(I)$ is the embedding of the image before finetuning on source, and $f_{T}(T)$ is the embedding of the text.
We verify the Alignment Hypothesis similarly to the Image Similarity Hypothesis, by binning samples using the AlignmentScore and computing accuracy for each bin using VL2V-SD. We provide the same analysis for many more DG methods in the Appendix Figure 8. In Figure 2 a.), we can see that the Alignment Hypothesis explains DG performance after source finetuning, significantly more strongly than for the Image Similarity Hypothesis in Figure 2 b.) This finding suggests that source fine-tuning in DG, which aims to achieve high performance across all target samples, only succeeds on those with high initial alignment.
4 Re-thinking Domain Generalization Benchmarking using the Alignment Hypothesis
Knowing that the Alignment Hypothesis holds for contrastively trained image-text models (Section 3), we can now use it as a tool to probe the performance of DG methods across different levels of pre-training alignment. We apply this approach to five widely-used DomainBed (Gulrajani & Lopez-Paz, 2020) DG datasets: VLCS (Fang et al., 2013), PACS (Li et al., 2017), OfficeHome (Ganin et al., 2016), TerraIncognita (Beery et al., 2018), and DomainNet  (Peng et al., 2019). This section discusses how we create new splits for existing DG datasets using AlignmentScore. We start by computing AlignmentScore for all samples in 5 DG datasets (Figure 3). Based on our observation that some samples are mislabeled, we perform dataset cleaning (4.1). Finally, we find an AlignmentScore threshold to split DG datasets into In-Pretraining (IP) and Out-of-Pretraining (OOP) evaluation subsets (Section 4.2), which we later use for evaluating DG methods (Section 5). In order to connect the AlignmentScore with the DG method, we use the same backbone both for splitting the datasets into IP and OOP subsets and for training DG methods.
Figure 3: Representative DomainBed dataset samples and their labels at various AlignmentScore values. At very low AlignmentScores, most labels (red boxes) are incorrect. At very high AlignmentScores, text present in the image corresponds to the label.
4.1 Data Exploration and Cleaning
We start by visualizing the data of all the datasets at various AlignmentScore values. We show some representative samples in Figure 3. At very low scores, we find that a large fraction of labels are incorrect (red boxes in Figure 3).
As a result, we divide the data into AlignmentScore intervals (e.g., 0.00-0.05, 0.05-0.10, and so on, up to 0.2) and randomly sample 100 instances from each interval for every dataset. This allows us to systematically analyze the relationship between AlignmentScore and label accuracy across different score ranges. For each interval, we then count the fraction of mislabelled samples to better understand how low AlignmentScores are associated with labeling errors. We find that below an AlignmentScore of 0.15, label noise is unacceptably high, with all datasets suffering the most from mislabelling (Table 4 in Appendix). Therefore, we discard all samples with AlignmentScore less than 0.15 in DomainBed-IP/OOP. As shown in Table 1, we observe that the percentage of discarded samples due to mislabeling varies across datasets, with VLCS and DomainNet having the highest rates at 12.41% and 7.64%, respectively.
Furthermore, on the right side of Figure 3, we observe that at very high AlignmentScores (greater than 0.4), the images often contain text directly related to the label. Since our goal is to evaluate visual recognition rather than text recognition (OCR), and CLIP is known to have strong OCR abilities  (Fort, 2021), we exclude all samples above AlignmentScore of 0.4 from DomainBed-IP/OOP. As shown in Table 1, although only a small portion of data is removed due to OCR filtering (0.00-0.15% across datasets), this issue may become more significant in future studies.
Table 1: Percentage of discarded samples due to mislabeling or label text in the image (OCR).
Dataset Dropped - Mislabelled Dropped - OCR
OfficeHome 0.92% 0.15%
PACS 3.05% 0.00%
TerraIncognita 0.22% 0.00%
DomainNet 7.64% 0.03%
VLCS 12.41% 0.00%
4.2 Data Splitting
After filtering, we focus on determining a threshold to split the dataset into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets. We select 0.21 as the threshold, based on the trends observed in Figure 2 (a.)), as this is the point where performance begins to improve significantly, indicating that existing methods become more effective. While this threshold represents a somewhat subjective choice informed by observed patterns, we provide AlignmentScores in the released data, allowing researchers the flexibility to experiment with their own thresholds.
Figure 4 illustrates how this split impacts the size and composition of each dataset. Certain domains naturally fall into the IP subset. For example, in DomainNet, the clipart domain is predominantly categorized as IP, likely due to its frequent presence on the internet and therefore the web-scraped pre-training data. On the other hand, TerraIncognita-OOP has relatively more balanced domains, but exhibits substantial class shift between OP and OOP splits (Figure 5), meaning some classes are better aligned than others during pre-training. More dataset statistics are in the Appendix.
Figure 4: DomainBed-(IP/OOP) Statistics: Breakdown of DomainBed-IP and DomainBed-OOP counts, by dataset and domain. Overall, DomainNet and VLCS have the largest fraction of samples falling into DomainBed-OOP.
Figure 5: Class-distribution shift: TerraIncognita’s class distribution differs between DB-IP and DB-OOP, indicating that some classes were better aligned during pretraining.
5 Experiments
5.1 Training and Evaluation Protocol
We aim to evaluate how well Domain Generalization (DG) methods perform across both pretraining-aligned(IP) and pretraining-misaligned(OOP) data. We adhere to the DomainBed evaluation methodology, where one domain is chosen as the target, and the remaining domains act as source domains. To maintain a sufficient amount of training data, we train all DG methods on the original, unsplit datasets. We use hyper-parameter values recommended by the original implementation authors for each method.
After training, models are evaluated separately on the IP and OOP subsets, as well as the original, unsplit test domain. This approach allows us to measure how well each method generalizes to both pretraining-aligned and pretraining-misaligned data.
We follow the literature’s standard practice of computing performance per target domain and averaging the results across all domains. This method ensures that any domain imbalances do not disproportionately influence the final performance metrics.
Table 2: Benchmarking DG methods on DB-IP/OOP
DomainBed-IP DomainNet OfficeHome PACS TI VLCS Average
CORAL (Sun & Saenko, 2016) 63.3 76.1 84.3 42.9 86.5 70.6
SAGM (Wang et al., 2023) 64.3 79.5 90.1 44.0 88.0 73.2
ERM* (Gulrajani & Lopez-Paz, 2020) 63.1 78.1 87.1 42.0 85.3 71.1
LP-FT (Kumar et al., 2021) 64.4 78.5 90.3 40.9 86.0 72.0
SWAD (Cha et al., 2021) 72.3 84.8 94.6 52.7 88.5 78.6
MIRO (Cha et al., 2022) 72.4 88.8 97.6 58.9 91.0 81.7
VL2V-SD (Addepalli et al., 2024) 78.1 91.4 98.0 48.1 92.4 81.6
Attn Tune (Teterwak et al., 2023) 69.2 84.8 96.4 53.0 88.7 78.4
MPA (Arpit et al., 2022) 73.6 85.1 95.4 54.4 90.7 79.8
CLIPOOD (Shu et al., 2023) 78.9 90.9 97.7 63.5 92.5 84.7
MIRO + SWAD 77.0 90.5 97.6 62.1 91.1 83.6
MIRO + MPA 78.2 90.7 98.1 62.6 91.0 84.1
Upper Bound (Target Finetune) 81.6 88.5 97.8 93.4 93.8 91.0
(a) Samples with high AlignmentScore values, indicating good alignment between image and text features learned during pretraining.
DomainBed-OOP DomainNet OfficeHome PACS TI VLCS Average
CORAL (Sun & Saenko, 2016) 22.3 42.6 74.1 16.0 74.0 45.8
SAGM (Wang et al., 2023) 23.0 44.5 74.2 19.3 73.3 46.9
ERM* (Gulrajani & Lopez-Paz, 2020) 22.3 42.9 76.9 16.5 76.4 47.0
LP-FT (Kumar et al., 2021) 22.7 43.4 78.6 23.1 70.7 47.7
SWAD (Cha et al., 2021) 28.6 49.9 79.1 21.0 77.0 51.1
MIRO (Cha et al., 2022) 28.4 56.6 84.7 18.5 73.7 52.4
VL2V-SD (Addepalli et al., 2024) 31.8 56.6 85.0 15.9 79.1 53.7
Attn Tune (Teterwak et al., 2023) 26.8 51.4 84.2 20.3 76.1 51.8
MPA (Arpit et al., 2022) 29.6 51.0 82.7 22.2 79.5 53.0
CLIPOOD (Shu et al., 2023) 33.9 63.9 87.2 19.9 80.7 57.1
MIRO + SWAD 32.0 59.0 85.4 21.1 78.9 55.3
MIRO + MPA 33.1 60.0 87.8 24.9 80.3 57.2
Upper Bound (Target Finetune) 48.8 61.9 92.9 83.2 92.4 75.8
(b) Samples with lower AlignmentScore values, representing cases where pretraining alignment is weak.
DomainBed-All DomainNet OfficeHome PACS TI VLCS Average
CORAL (Sun & Saenko, 2016) 50.6 73.2 83.2 39.6 78.5 65.0
SAGM (Wang et al., 2019) 51.5 76.4 87.5 41.0 80.4 67.3
ERM* (Gulrajani & Lopez-Paz, 2020) 50.5 75.0 85.2 39.0 77.9 65.5
LP-FT (Kumar et al., 2021) 51.3 75.5 88.4 38.5 78.0 66.3
SWAD (Cha et al., 2021) 57.9 81.8 92.4 49.0 80.1 72.2
MIRO (Cha et al., 2022) 57.5 85.8 96.4 54.3 81.1 75.0
VL2V-SD (Addepalli et al., 2024) 62.0 88.3 96.9 44.4 82.7 74.9
Attn Tune (Teterwak et al., 2023) 55.4 81.9 95.4 49.1 81.8 72.7
MPA (Arpit et al., 2022) 58.9 82.0 94.3 50.7 82.3 73.6
CLIPOOD (Shu et al., 2023) 63.6 88.3 96.8 58.5 83.4 78.1
MIRO + SWAD 61.4 87.6 96.6 57.4 82.0 77.0
MIRO + MPA 62.4 87.9 97.2 58.2 82.8 77.7
Upper Bound (Target Finetune) 70.4 86.2 97.2 92.4 87.9 86.8
(c) Performance of DG methods on unsplit DomainBed
5.2 Algorithms
Explicit Regularization towards Pretraining: Several recent DG methods leverage explicit regularization towards the initialization. These methods generally operate either in weight space (by regularizing or freezing model parameters) or in feature space (by aligning internal feature representations with those of the pretrained model).
  • MIRO (Cha et al., 2022): Minimizes the Mutual Information between DG model intermediate features and CLIP intermediate features.
  • Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022): Freezes all parameters except those in the Multiheaded-Attention Layers.
  • VL2V-SD (Addepalli et al., 2024): Self-distills a linear combination of CLIP vision and text outputs into a model.
  • CLIPood (Shu et al., 2023): Regularizes both weights and output features. Weights are averaged between the pre-trained CLIP model and the fine-tuned model, and outputs are regularized using a loss that incorporates class information from the pre-trained text encoder.
  • Linear Probe - Fine Tuning (LP-FT) (Kumar et al., 2021): Freezes the backbone, trains a linear probe, and then performs full finetuning. An untrained linear probe can cause finetuning to update the frozen backbone needlessly, potentially unlearning discriminative features. This biases the model towards pre-trained weights with smaller gradient updates.
Domain Invariance: A classic idea for Domain Generalization is Domain Invariance in the feature space, where the model learns only class-discriminative features shared among all training/source domains. CORAL (Sun & Saenko, 2016) matches the second moments of features across different domains and has been shown to be highly effective (Gulrajani & Lopez-Paz, 2020).
Flat Optima: Studies by Izmailov et al. (2018) and Cha et al. (2021) have shown that flat minima generalize better than sharp minima, as they make loss values less sensitive to perturbations in the loss surface, resulting in smaller increases in loss during domain shifts.
  • SWAD (Cha et al., 2021): Averages model parameters during training, determining the interval over which to average using validation loss over source domain.
  • Model Parameter Averaging (MPA) (Arpit et al., 2022): Starts averaging model parameters after a number of burn-in steps to find flat minima.
  • SAGM (Wang et al., 2023): An optimizer that explicitly optimizes for flat minima.
Baseline and Oracle: We also train a baseline and oracle model for lower-bound and upper-bound reference. The baseline model is an Empirical Risk Minimization (ERM) model that is finetuned on source domains and evaluated on target domains, and has been found to be effective for the DG task (Gulrajani & Lopez-Paz, 2020). The oracle model is trained on an 80% training split of all domains and evaluated on a 20% test split. The oracle model removes the OOD aspect of generalization and provides a reasonable upper bound for DG methods.
5.3 Results
We show results in Table 2(c) for DomainBed-OOP, DomainBed-IP, and standard DomainBed datasets. Underlined results represent the best performance of any single method (excluding method combinations), while the bold numbers show the highest performance overall, excluding the upper bound.
DG methods have nearly solved DomainBed-IP: In most datasets within DomainBed-IP, domain generalization approaches have nearly solved the problem of generalization. On three out of five datasets (DomainNet, OfficeHome, and PACS), the best DG method even outperforms the oracle! A notable exception is TerraIncognita, where CLIPood scores only 63.5%, far below the oracle’s 93.4%, highlighting this dataset as a challenging outlier. Interestingly, all three datasets where performance exceeds the oracle have an average IP AlignmentScore of 0.28, while the others have a lower average AlignmentScore of around 0.26. Therefore, the underperformance of TerraIncognita may be partially explained by its lower IP AlignmentScore, suggesting that alignment plays a significant role in DG performance, even within the IP case.
DG Methods leave much to be desired on DomainBed-OOP, but are still stronger than ERM: In DomainBed-OOP, we observe that even the top-performing DG methods struggle with low-alignment data. For example, CLIPood achieves 57.1% accuracy, which is a significant drop compared to its performance on DomainBed-IP (84.7%) and DomainBed-All (78.1%). Despite this, DG methods still outperform Empirical Risk Minimization (ERM), which scores only 47.0% on DomainBed-OOP. Therefore DG methods are better equipped to handle domain shifts than ERM, possibly due to weak transfer of knowlegdge from pre-trained features. Nevertheless, there is still substantial room for improvement on low-alignment samples.
State-of-the-Art methods do not consistently outperform older methods on OOP data: While CLIPood (Shu et al., 2023) clearly outperforms other methods on DomainBed-All with 78.1% accuracy, it performs comparably to older methods on DomainBed-OOP. For instance, MIRO + MPA (Cha et al., 2022; Arpit et al., 2022) achieves 57.2% on DomainBed-OOP, which is nearly the same as CLIPood’s 57.1%. This suggests that CLIPood’s primary advantage comes from well-aligned samples in DomainBed-IP, where it reaches 84.7% accuracy, half a percent better than the next-best method.
Model Parameter Averaging (MPA) boosts performance on OOP data: MPA performs significantly better than ERM on DomainBed-OOP, achieving 53.0% compared to ERM’s 47.0%. This is half a percent better than MIRO on OOP data, despite being 2% worse than MIRO on IP data. When combined with MIRO, MPA delivers the best performance on DomainBed-OOP, scoring 57.2%, slightly surpassing CLIPood. This suggests that MPA can complement other regularization-based methods like MIRO. On DomainBed-IP, MIRO + MPA scores 84.1%, close to CLIPood’s 84.7%, demonstrating versatility across both high- and low-alignment data. Interestingly, SWAD underperforms MPA on DomainBed-OOP by 2%, despite being conceptually similar. We attribute this to selecting the averaging interval on the source data, which introduces overfitting to source domains.
6 Conclusion
In this paper, we systematically explore how Domain Generalization (DG) methods rely on pre-trained feature alignment from models like CLIP. We hypothesize that the alignment between image and text embeddings during pre-training strongly predicts DG performance. Our experiments confirm this, showing that methods perform well on high-alignment samples (DomainBed-IP) but struggle on low-alignment data (DomainBed-OOP). While we focus on multi-source DG, we expect similar results in single-source finetuning due to the foundational role of pre-trained alignment. Notably, state-of-the-art methods like CLIPood perform near oracle-level on aligned data but see significant drops on misaligned samples. This suggests current DG methods rely on pre-trained features and fail to learn new, generalizable features from source domains. Moving forward, two paths emerge: developing DG methods that better learn generalizable features, or focusing on improving pre-trained backbones. While foundation models will continue to advance, there will always be specialized distributions where they fail. We hope these findings inspire further research into improving generalization on low-alignment data, pushing DG beyond reliance on pre-trained alignment.
DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited. This material is based upon work supported by the Under Secretary of Defense for Research and Engineering under Air Force Contract No. FA8702-15-D-0001. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Under Secretary of Defense for Research and Engineering.
Appendix A Appendix
A.1 Training and Evaluation Details
Code is zipped in the supplementary.
We use a slightly modified MIRO Cha et al. (2022) codebase for training and evaluation. We use leave-one-out evaluation, where a model is trained on all domains except the evaluation domain. We emphasize that we use DomainBed-IP and DomainBed-OOP as evaluation data only, models are trained on full datasets.
For training, we use an OpenCLIP-ViT-B/16 Ilharco et al. (2021) trained on LAION-400M Schuhmann et al. (2021). We use default hyper-parameters as defined by  Cha et al. (2022). This includes a learning rate of 5e-5, weight decay of 0.0, a batch size of 32 per-domain, an Adam Optimizer, and no dropout for all methods.
For evaluation, unlike DomainBed, we consider the entire test domain instead of an 80% random split. Following standard practice, we first compute accuracy for each domain, then average those accuracies to get dataset level statistics, and finally compute overall averages averaging across datasets.
For benchmarked methods, we also use hyper-parameters found to be best in respective papers. For SWAD, we use an optimum patience parameter value of 3, overfit patience parameter value of 6, and tolerance ratio of 6. For MIRO, we use use regularizer loss weight of 1.0. For CORAL, we use a CORAL regularizer weight of 1.0, following  Cha et al. (2021). For LP-FT, we train the linear probe for 600 steps before unlocking the full backbone. For Model Parameter Averaging, we burn in the training for 600 steps before averaging iterates. For VL2V-SD and CLIPood, we directly use the author’s implementation and hyper-parameters, except initializing with OpenCLIP Ilharco et al. (2021).
A.2 Training Compute
Each run uses an A6000 48GB GPU, trained for up to 12 hours per domain-dataset combination.
Appendix B Additional Results
B.1 AlignmentScore vs Accuracy
In Figure 8, we plot all benchmarked methods from the main paper, with x-axis corresponding to AlignmentScore, and the y-axis corresponding to the Top-1 Accuracy. We normalize for dataset size, so that no dataset dominates the count. In Figures 9 through 13, we plot these statistics independently per dataset, and find the trends consistent across datasets.
B.2 Per-Dataset Benchmarking Results
We expand Table 2(c) in the main paper into per-dataset results in Table 5 through 19.
B.3 Similarity of Target to Pre-training
To evaluate the Image Similarity Hypothesis, we retrieve the nearest neighbors from the Laion-400M dataset Schuhmann et al. (2021). This raises the question of how similar the target domains are to the pre-training data and whether the source domains might be even more similar. To investigate this, we compute Maximum Mean Discrepancy (MMD) distances between PACS domains and their nearest neighbors from Laion-400M, as shown in Figure 7. Our results indicate that target domains are, in fact, more similar to the pre-training data than source domains. We inspect nearest neighbors manually, and find even exact duplicates (Figure 6). Interestingly, while we found not only domain-level duplicates but also exact matches in the pre-training data, the Image Similarity Hypothesis is ultimately less predictive than the Alignment Hypothesis.
Figure 6: Nearest neighbors of target images in pre-training LAION data.
Figure 7: MMD between pre-training and target, source and target, and target and target for PACS. Target is more similar to pre-training than source. Despite this, Alignment is a better predictor of DG performance than perceptual similarity.
B.4 Other backbones
We benchmark 2 additional backbones(DINOv2 (Oquab et al., 2023) and OpenAI CLIP) using the MIRO + MPA Domain Generalization method, which we found to be the strongest in our paper, on two datasets (OfficeHome and PACS). This consistency is likely due to the similar nature of the pre-training datasets, both sourced from web scraping and of comparable scale. These findings reinforce that the usefulness of the DomainBed-IP/OOP split is not confined to a specific backbone.
Backbone OfficeHome - IP PACS - IP OfficeHome - OOP PACS - OOP
OpenCLIP-ViT/B 16 90.7 98.1 60.0 87.8
CLIP-ViT/B 16 88.1 97.6 57.0 86.9
DinoV2 87.4 97.1 58.9 85.0
Table 3: Benchmark results of different backbones on OfficeHome and PACS datasets.
Appendix C Additional DomainBed-(IP/OOP) Statistics and Analysis
C.1 Constituent Datasets
We split 5 datasets in DomainBed-(IP/OOP),( VLCS Fang et al. (2013), DomainNet Peng et al. (2019), OfficeHome Ganin et al. (2016), PACS Li et al. (2017), and TerraIncognita Beery et al. (2018)). Here we provide basic statistics of each.
VLCS has 5 classes and 4 domains: Caltech101, LabelME, SUN09, and VOC2007, with 10729 samples. The domain shift is dataset source.
DomainNet contrains 345 classes and 6 domains: clipart, infograph, quickdraw, real, and sketch. It has a total of 586,575 samples. The dataset shift is style.
OfficeHome has 65 classes and 4 domains: art, clipart, product, and real. The dataset shift is style.
TerraIncognita has 10 classes of wildlife cameras. There are 4 domains of different cameras and 24788 samples. The dataset shift is camera location.
PACS has 9991 sampes and 4 domains: arts, cartoon, photo, and sketch. There are 7 classes. The dataset shift is style.
C.2 Mislabeling Rates
In Table 4 we present the mislabelling rates at various AlingmentScore values.
Table 4: Mislabeling rates across different AlignmentScore ranges.
Dataset 0.0-0.05 0.05-0.10 0.10-0.15 0.15-0.20
OfficeHome 100% 45% 28% 12%
PACS - 46% 5% 0%
TerraIncognita - 55% 34% 23%
DomainNet 65% 35% 33% 11%
VLCS 33% 14% 9% 3%
C.3 Class Distribution of DomainBed-(IP/OOP):
In Figures 14 through 28, we provide class distribution statistics of different datasets before splitting and in our IP and OOP splits. We find some interesting patterns. For example, in Office-Home, the OOP class is dominated by marker and toys, while the IP split has a much more uniform distribution. Similarly, both PACS (DomainBed-OOP) and VLCS (DomainBed-OOP) are dominated by person.
Figure 8: Plotting AlignmentScore vs Top-1 Accuracy for all benchmarked methods, for all datasets together. Although some methods are stronger than others, all follow the same trend of increasing accuracy with increased AlignmentScore.
Figure 9: Plotting AlignmentScore vs Top-1 Accuracy on the PACS dataset.
Figure 10: Plotting AlignmentScore vs Top-1 Accuracy on the DomainNet dataset.
Figure 11: Plotting AlignmentScore vs Top-1 Accuracy on the OfficeHome dataset.
Figure 12: Plotting AlignmentScore vs Top-1 Accuracy on the TerraIncognita dataset.
Figure 13: Plotting AlignmentScore vs Top-1 Accuracy on the VLCS dataset.
Table 5: Per-domain breakdown for OfficeHome (DomainBed-OOP)
Method Art Clipart Product Real Avg
CORALSun & Saenko (2016) 33.0 29.4 50.0 58.0 42.6
SAGMWang et al. (2023) 36.8 30.7 51.7 58.7 44.5
ERM*Gulrajani & Lopez-Paz (2020) 36.0 28.1 51.3 56.0 42.9
LP-FTKumar et al. (2021) 37.2 29.7 52.0 54.7 43.4
SWADCha et al. (2021) 43.9 37.1 56.0 62.7 49.9
MIRO Cha et al. (2022) 49.4 39.6 62.7 74.7 56.6
VL2V-SD Addepalli et al. (2024) 56.4 39.3 65.3 65.3 56.6
Attn. Tune Teterwak et al. (2023); Touvron et al. (2022) 46.0 36.1 59.7 64.0 51.4
Model Parameter Averaging (MPA) Arpit et al. (2022) 46.0 36.4 55.7 66.0 51.0
CLIPOOD (Shu et al., 2023) 60.3 44.4 73.3 78.0 63.9
MIRO + SWAD 55.2 44.7 63.3 72.7 59.0
MIRO + MPA 53.5 44.1 65.7 76.7 60.0
Table 6: Per-domain breakdown for OfficeHome (DomainBed-IP)
Method Art Clipart Product Real Avg
CORAL 71.6 62.9 85.8 84.0 76.1
SAGM 75.7 70.8 87.0 84.5 79.5
ERM* 74.1 67.4 86.7 84.1 78.1
LP-FT 75.2 66.9 88.9 83.0 78.5
SWAD 81.4 77.8 91.8 88.2 84.8
MIRO 88.3 81.1 93.7 92.1 88.8
VL2V-SD 91.0 83.8 96.7 94.2 91.4
Attn. Tune 86.0 75.1 89.5 88.7 84.8
Model Parameter Averaging (MPA) 83.6 77.1 91.1 88.5 85.1
CLIPOOD (Shu et al., 2023) 92.3 81.3 96.0 94.2 90.9
MIRO + SWAD 90.4 83.2 95.5 92.8 90.5
MIRO + MPA 90.6 83.9 95.4 92.9 90.7
Table 7: Per-domain breakdown for OfficeHome (DomainBed-All)
Method Art Clipart Product Real Avg
CORAL 66.6 60.1 83.0 83.0 73.2
SAGM 70.7 67.1 84.3 83.6 76.4
ERM* 69.2 63.9 83.9 83.1 75.0
LP-FT 70.3 63.6 86.0 82.0 75.5
SWAD 76.6 74.3 89.0 87.2 81.8
MIRO 83.2 77.5 91.2 91.4 85.8
VL2V-SD 86.4 79.7 94.1 93.1 88.3
Attn. Tune 80.8 71.7 87.2 87.8 81.9
Model Parameter Averaging (MPA) 78.7 73.4 88.4 87.6 82.0
CLIPOOD (Shu et al., 2023) 87.9 77.8 94.1 93.6 88.3
MIRO + SWAD 85.6 79.9 93.0 92.1 87.6
MIRO + MPA 85.6 80.5 93.0 92.3 87.9
Table 8: Per-domain breakdown for TerraIncognita (DomainBed-OOP)
Method L100 L38 L43 L46 Avg
CORAL 13.9 8.1 21.1 21.1 16.0
SAGM 17.2 7.8 20.2 32.2 19.3
ERM* 13.1 9.6 22.9 20.5 16.5
LP-FT 16.7 5.1 25.6 44.9 23.1
SWAD 20.7 8.2 26.9 28.2 21.0
MIRO 19.4 6.6 30.0 17.9 18.5
VL2V-SD 9.3 5.6 20.2 28.4 15.9
Attn. Tune 20.5 10.7 24.8 25.3 20.3
Model Parameter Averaging (MPA) 21.7 8.6 30.9 27.6 22.2
CLIPOOD (Shu et al., 2023) 37.9 11.3 21.6 8.9 19.9
MIRO + SWAD 22.5 5.9 30.3 25.9 21.1
MIRO + MPA 24.8 9.6 32.5 32.8 24.9
Table 9: Per-domain breakdown for TerraIncognita (DomainBed-IP)
Method L100 L38 L43 L46 Avg
CORAL 46.4 35.3 56.9 32.8 42.9
SAGM 55.1 38.7 52.5 29.5 44.0
ERM* 37.4 38.0 55.2 37.5 42.0
LP-FT 50.7 36.5 58.1 18.2 40.9
SWAD 60.1 41.7 65.7 43.5 52.7
MIRO 69.5 50.2 69.7 46.1 58.9
VL2V-SD 52.8 38.4 57.6 43.8 48.1
Attn. Tune 52.4 44.0 68.0 47.5 53.0
Model Parameter Averaging (MPA) 63.8 41.4 67.9 44.7 54.4
CLIPOOD (Shu et al., 2023) 77.7 56.4 69.0 51.0 63.5
MIRO + SWAD 68.9 54.8 73.4 51.2 62.1
MIRO + MPA 69.4 55.9 73.6 51.4 62.6
Table 10: Per-domain breakdown for TerraIncognita (DomainBed-All)
Method L100 L38 L43 L46 Avg
CORAL 43.6 32.7 50.4 31.8 39.6
SAGM 51.9 35.8 46.6 29.7 41.0
ERM* 35.3 35.3 49.3 36.0 39.0
LP-FT 47.8 33.5 52.1 20.5 38.5
SWAD 56.7 38.5 58.6 42.2 49.0
MIRO 65.2 46.0 62.4 43.7 54.3
VL2V-SD 49.1 35.3 50.8 42.5 44.4
Attn. Tune 49.7 40.8 60.1 45.6 49.1
Model Parameter Averaging (MPA) 60.1 38.3 61.1 43.2 50.7
CLIPOOD (Shu et al., 2023) 74.3 52.1 60.4 47.4 58.5
MIRO + SWAD 64.9 50.1 65.5 49.0 57.4
MIRO + MPA 65.6 51.5 66.0 49.8 58.2
Table 11: Comparison of Methods for DomainNet Dataset (DomainBed-OOP)
Method Clp Inf Pnt Qkdr Real Skt Avg
CORAL 34.9 19.2 26.9 5.7 24.6 22.8 22.3
SAGM 36.4 19.4 25.1 6.1 26.8 24.2 23.0
ERM* 36.2 19.0 23.4 5.4 26.6 23.4 22.3
LP-FT 34.7 20.9 25.2 6.1 25.9 23.2 22.7
SWAD 42.5 28.3 33.4 7.6 30.2 29.5 28.6
MIRO 42.5 37.0 27.9 3.5 33.5 25.8 28.4
VL2V-SD 43.0 43.2 34.8 3.0 37.1 29.5 31.8
Attn. Tune 40.4 29.1 29.2 4.3 31.3 26.6 26.8
Model Parameter Averaging (MPA) 42.3 30.9 34.9 7.4 31.6 30.7 29.6
CLIPOOD (Shu et al., 2023) 39.2 50.2 39.5 2.1 40.9 31.3 33.9
MIRO + SWAD 46.1 40.4 34.0 4.4 35.2 31.7 32.0
MIRO + MPA 46.2 42.6 36.0 4.4 36.0 33.3 33.1
Table 12: Comparison of Methods for DomainNet Dataset (DomainBed-IP)
Method Clp Inf Pnt Qkdr Real Skt Avg
CORAL 80.3 52.3 71.1 33.2 71.7 71.2 63.3
SAGM 81.7 52.8 70.2 34.5 73.4 73.1 64.3
ERM* 80.6 53.4 69.0 31.3 73.0 71.6 63.1
LP-FT 80.0 56.5 70.5 34.0 73.9 71.5 64.4
SWAD 85.5 67.1 80.2 41.3 80.1 79.8 72.3
MIRO 85.4 77.7 79.1 34.6 83.2 74.2 72.4
VL2V-SD 88.3 85.6 84.6 38.0 88.3 84.0 78.1
Attn. Tune 83.5 66.1 77.3 32.6 80.3 75.5 69.2
Model Parameter Averaging (MPA) 85.8 69.7 81.6 41.4 81.4 81.4 73.6
CLIPOOD (Shu et al., 2023) 86.6 85.4 87.6 38.0 91.0 84.7 78.9
MIRO + SWAD 88.2 81.7 83.9 40.8 85.8 81.9 77.0
MIRO + MPA 88.1 83.7 85.5 41.3 86.5 84.0 78.2
Table 13: Comparison of Methods for DomainNet Dataset (DomainBed-All)
Method Clp Inf Pnt Qkdr Real Skt Avg
CORAL 72.7 27.1 58.0 20.1 67.1 58.8 50.6
SAGM 74.1 27.4 56.9 20.9 68.9 60.6 51.5
ERM* 73.1 27.4 55.7 19.0 68.5 59.3 50.5
LP-FT 72.4 29.3 57.1 20.7 69.2 59.2 51.3
SWAD 78.0 36.0 66.1 25.2 75.2 66.7 57.9
MIRO 78.0 42.9 64.0 20.0 78.3 61.7 57.5
VL2V-SD 80.5 47.8 69.5 21.6 83.1 69.8 62.0
Attn. Tune 76.1 36.0 63.0 19.2 75.4 62.9 55.4
Model Parameter Averaging (MPA) 78.3 37.9 67.4 25.2 76.5 68.2 58.9
CLIPOOD (Shu et al., 2023) 78.4 52.9 72.6 21.2 85.9 70.7 63.6
MIRO + SWAD 80.8 45.6 68.8 23.7 80.7 68.7 61.4
MIRO + MPA 80.7 47.2 70.4 23.9 81.4 70.6 62.4
Table 14: Per-domain breakdown for VLCS (DomainBed-OOP)
Method Caltech101 LabelMe SUN09 VOC2007 Avg
CORAL 93.3 54.6 72.9 75.3 74.0
SAGM 80.0 58.8 76.5 78.0 73.3
ERM* 100.0 57.1 73.3 75.3 76.4
LP-FT 80.0 55.4 71.8 75.7 70.7
SWAD 100.0 51.2 79.1 77.7 77.0
MIRO 86.7 52.6 82.9 72.7 73.7
VL2V-SD 100.0 55.0 80.4 81.1 79.1
Attn. Tune 86.7 61.0 77.9 79.0 76.1
Model Parameter Averaging (MPA) 100.0 53.6 83.6 80.9 79.5
CLIPOOD (Shu et al., 2023) 100.0 55.0 84.9 82.8 80.7
MIRO + SWAD 100.0 56.7 84.7 79.9 80.3
MIRO + MPA 100.0 54.6 83.4 77.4 78.9
Table 15: Per-domain breakdown for VLCS (DomainBed-IP)
Method Caltech101 LabelMe SUN09 VOC2007 Avg
CORAL 98.2 83.5 84.5 79.8 86.5
SAGM 95.9 84.3 86.3 85.5 88.0
ERM* 96.9 83.7 86.1 74.6 85.3
LP-FT 97.1 83.2 82.8 80.7 86.0
SWAD 97.9 82.9 88.4 84.8 88.5
MIRO 98.9 83.3 93.5 88.2 91.0
VL2V-SD 99.5 83.9 91.8 94.6 92.4
Attn. Tune 98.0 84.7 85.7 86.5 88.7
Model Parameter Averaging (MPA) 98.4 83.0 92.6 88.8 90.7
CLIPOOD (Shu et al., 2023) 98.9 83.6 93.5 94.0 92.5
MIRO + SWAD 97.6 83.4 92.6 90.6 91.1
MIRO + MPA 98.0 83.2 92.9 90.1 91.0
Table 16: Per-domain breakdown for VLCS (DomainBed-All)
Method Caltech101 LabelMe SUN09 VOC2007 Avg
CORAL 98.2 66.2 72.7 77.0 78.5
SAGM 95.8 68.8 76.2 80.8 80.4
ERM* 96.9 67.8 73.0 73.9 77.9
LP-FT 97.0 66.4 71.4 77.4 78.0
SWAD 98.0 64.5 77.7 80.3 80.1
MIRO 98.7 65.3 81.2 79.3 81.1
VL2V-SD 99.5 66.6 78.8 86.0 82.7
Attn. Tune 97.9 70.3 77.2 81.9 81.8
Model Parameter Averaging (MPA) 98.4 65.7 81.3 83.7 82.3
CLIPOOD (Shu et al., 2023) 98.9 66.7 81.6 86.6 83.4
MIRO + SWAD 97.6 66.4 81.6 82.6 82.0
MIRO + MPA 98.0 67.4 82.2 83.6 82.8
Table 17: Per-domain breakdown for PACS (DomainBed-OOP)
Method Art Cartoon Photo Sketch Avg
CORAL 80.4 84.3 89.5 42.2 74.1
SAGM 74.4 78.1 89.5 54.7 74.2
ERM* 75.9 86.8 91.6 53.2 76.9
LP-FT 74.2 85.9 97.9 56.3 78.6
SWAD 84.6 81.6 95.8 54.3 79.1
MIRO 94.2 94.1 97.9 52.8 84.7
VL2V-SD 93.5 96.5 97.9 52.2 85.0
Attn. Tune 94.0 91.4 95.8 55.7 84.2
Model Parameter Averaging (MPA) 94.0 88.9 97.9 50.2 82.7
CLIPOOD (Shu et al., 2023) 96.0 96.8 97.9 58.3 87.2
MIRO + SWAD 96.0 94.9 100.0 50.7 85.4
MIRO + MPA 95.3 95.4 100.0 60.3 87.8
Table 18: Per-domain breakdown for PACS (DomainBed-IP)
Method Art Cartoon Photo Sketch Avg
CORAL 83.3 89.8 91.8 72.1 84.3
SAGM 89.7 93.0 94.7 82.9 90.1
ERM* 88.7 93.0 91.5 75.2 87.1
LP-FT 85.3 93.7 97.9 84.2 90.3
SWAD 96.4 93.0 97.8 91.0 94.6
MIRO 99.2 96.7 99.9 94.4 97.6
VL2V-SD 99.6 98.9 99.9 93.6 98.0
Attn. Tune 97.3 97.6 98.9 91.7 96.4
Model Parameter Averaging (MPA) 98.6 95.8 98.2 89.2 95.4
CLIPOOD (Shu et al., 2023) 99.4 99.3 99.9 92.0 97.7
MIRO + SWAD 99.1 97.9 99.9 93.3 97.6
MIRO + MPA 99.0 98.5 100.0 94.7 98.1
Table 19: Per-domain breakdown for PACS (DomainBed-All)
Method Art Cartoon Photo Sketch Avg
CORAL 81.4 89.2 91.7 70.5 83.2
SAGM 83.4 90.4 94.6 81.4 87.5
ERM* 83.0 92.1 91.5 74.0 85.2
LP-FT 80.5 92.7 97.9 82.7 88.4
SWAD 91.5 91.2 97.8 89.0 92.4
MIRO 97.1 96.4 99.8 92.2 96.4
VL2V-SD 97.7 98.5 99.9 91.3 96.9
Attn. Tune 96.4 96.5 98.8 89.7 95.4
Model Parameter Averaging (MPA) 97.2 94.8 98.2 87.1 94.3
CLIPOOD (Shu et al., 2023) 98.2 98.8 99.8 90.2 96.8
MIRO + SWAD 98.0 97.4 99.9 91.0 96.6
MIRO + MPA 97.8 98.0 99.9 92.9 97.2
Figure 14: Class distribution of DomainNet-IP. Zoom in on pdf for best viewing.
Figure 15: Class distribution of DomainNet-OOP. Zoom in on pdf for best viewing.
Figure 16: Class distribution of DomainNet before splitting. Zoom in on PDF before viewing.
Figure 17: Class distribution of OfficeHome-IP. Zoom in on pdf for best viewing.
Figure 18: Class distribution of OfficeHome-OOP. Zoom in on pdf for best viewing.
Figure 19: Class distribution of OfficeHome before Splitting. Zoom in on pdf for best viewing.
Figure 20: Class distribution of PACS-IP.
Figure 21: Class distribution of PACS-OOP.
Figure 22: Class distribution of PACS before splitting.
Figure 23: Class distribution of TerraIncognita-IP
Figure 24: Class distribution of TerraIncognita-OOP
Figure 25: Class distribution of TerraIncognita before splitting.
Figure 26: Class distribution of VLCS-IP
Figure 27: Class distribution of VLCS-OOP
Figure 28: Class distribution of VLCS before splitting.

### v3

@@
 Put together, our findings highlight the need for DG methods which can generalize beyond pretraining alignment.
+We release DomainBed-OOP at https://huggingface.co/datasets/PTeterwak/DomainBed_OOP.
 1 Introduction
@@
 DG pipelines typically consist of three stages: pretraining a model on a large, general dataset; finetuning the model with one or more source domains; and finally evaluating the model on target domains that are distinct from source domains.
-More and more, DG methods rely on huge-scale foundation models for initialization.
+DG methods increasingly rely on huge-scale foundation models for initialization (e.g.(Shu et al., 2023; Cha et al., 2022; Addepalli et al., 2024)).
 Simultaneously, finetuning has increasingly incorporated regularization to prevent catastrophic forgetting.
-As a result, it remains unclear whether DG adaptation methods are genuinely improving or if enhanced benchmark performance is simply due to stronger pre-training combined with regularization, or even the presence of target domains within the hundred million-scale pre-training data.
-In this work, we examine the reliance of recent DG methods on pre-trained features.
-We focus on CLIP-based models, as they are used in all state-of-the-art DG methods (Addepalli et al., 2024; Cho et al., 2023; Shu et al., 2023; Mao et al., 2024).
-While prior studies  (Mayilvahanan et al., 2024) have shown that perceptual similarity to pre-training data explains zero-shot performance—referred to as the Image Similarity Hypothesis—we find this relationship to be limited in the DG setting.
-Despite evidence of target domains being present in pre-training (Figure 6), perceptual similarity alone does not fully explain accuracy in the DG context.
-We propose that it is not just the presence of similar data in pre-training that matters, but also how well this data was learned.
+As a result, it remains unclear whether DG methods are genuinely improving or if benchmark performance gains are simply due to stronger pre-training, possibly with the target domains within the hundred million-scale pre-training data (Mayilvahanan et al., 2025), combined with regularization.
+In this work, we examine the reliance of recent state-of-the-art CLIP-based DG methods on pre-trained features.
+While prior studies have shown that perceptual similarity to pre-training data explains zero-shot performance—referred to as the Image Similarity Hypothesis (Mayilvahanan et al., 2024)—we find this relationship to be limited in the DG setting.
+Despite evidence of target domains being present in pre-training (Figure 4), we find that perceptual similarity alone does not fully explain accuracy in the DG context (Section 3).
+We argue that it is not just the presence of similar data in pre-training that matters, but also how well this data was learned.
 To this end, we introduce the Alignment Hypothesis, which states that pre-trained alignment between image and class embeddings is still predictive of DG performance even after source finetuning.
-We find that, performance for low alignment samples can be almost 0, while performance for high alignment samples is close to perfect.
+We note that we do not make assumptions of how or why alignment arose.
+In the Alignment Hypothesis, pre-training alignmnent is used as a measure for how well a sample is learned.
+We find that performance for low alignment samples can be almost 0, while performance for high alignment samples is close to perfect.
 These results confirm the Alignment Hypothesis.
@@
 3) Our analysis shows that DG methods only result in strong alignment for IP data, leaving OOP data misaligned (Figure 2).
-The confirmation of the Alignment Hypothesis gives us a tool to separate in-pretraining (IP) data from out-of-pretraining (OOP) data for a particular backbone, and we do so for five DG datasets with OpenCLIP-ViT/B-16.
+The confirmation of the Alignment Hypothesis gives us a tool to separate aligned and well learned in-pretraining (IP) data from misaligned and poorly learned out-of-pretraining (OOP) data for a particular backbone, and we do so for five DG datasets with OpenCLIP-ViT/B-16.
 We call the resulting splits DomainBed-IP and DomainBed-OOP.
 Evaluating on DomainBed-IP/OOP offers a view of where current DG methods fail and where they succeed.
-We find that all methods, including those considered state-of-the-art, perform poorly on OOP data, i.e.
-data that the pretrained backbone hadn’t already aligned well.
-Furthermore, recent state-of-the-art methods do not outperform older methods on OOP data.
-For example, CLIPood  (Shu et al., 2023) slightly under performs a combination of older methods (MIRO (Cha et al., 2022) + MPA (Arpit et al., 2022)) on DomainBed-OOP.
+We focus on CLIP-based methods, as they are used in state-of-the-art DG methods (Addepalli et al., 2024; Cho et al., 2023; Shu et al., 2023; Mao et al., 2024); we believe extensions to pure vision models such as DINOv2 (Oquab et al., 2023) represent interesting future work.
+We find that all methods, including those considered state-of-the-art, perform poorly on OOP data, i.e., data that the pretrained backbone hadn’t already aligned well.
+Furthermore, recent state-of-the-art methods do not outperform older methods on OOP data; CLIPood  (Shu et al., 2023) slightly under performs a combination of older methods (MIRO (Cha et al., 2022) + MPA (Arpit et al., 2022)) on DomainBed-OOP.
 At the same time, existing DG methods show exceptional performance on DomainBed-IP, sometimes even outperforming an oracle model trained on the target domain.
-These results suggest that future research should aim to enhance DG methods on low-alignment data while preserving the already strong performance on high-alignment data.
+These results suggest that future research should aim to enhance DG methods on low-alignment data.
 In summary, we make the following contributions:
@@
   • Propose a new IP/OOP evaluation framework: We demonstrate that splitting target data by its alignment with the pre-trained backbone can effectively test Out-of-Pretraining (OOP) generalization.
-We will release IP/OOP splits for the DomainBed dataset to support future research.
+We release IP/OOP splits for DomainBed datasets.
   • Expose strengths and limitations of state-of-the-art DG methods: Using DomainBed-IP/OOP we find that leading DG methods perform well on data well-aligned by pre-training but struggle on misaligned samples, emphasizing the need for methods that move beyond reliance on pre-training.
@@
 Multi-Source Domain Generalization: Domain Generalization aims to mitigate the impacts of domain shifts between source (training) and target (deployment) domains.
+These can include subpopulation shifts where all testing subpopulations are present in the training data but in different proportions (Dehdashtian et al., 2024), or it could be the case we consider in this work where the testing subpopulation is not at all present in the training subpopulation.
+Although we focus on the multi-source domain generalization task, where domains between train and test are completely disjoint, our analysis can be extended to other types of generalization.
 One standard approach is domain-invariant feature learning, which leverages domain labels to learn domain-invariant features.
-CORAL  (Sun & Saenko, 2016) aligns second-order statistics, while DANN  (Ganin et al., 2016) uses an adversarial loss.
+CORAL (Sun & Saenko, 2016) aligns second-order statistics, while DANN (Ganin et al., 2016) and ADDA (Tzeng et al., 2017) uses an adversarial loss.
 Gulrajani & Lopez-Paz (2020) show that ERM, which does not align features between domains, can outperform most prior work while being easier to tune.
@@
 This works under the assumption that pretrained features are useful for target data, and should not be unlearned.
-The general idea can be applied to weight space (L2SP (Xuhong et al., 2018)), feature space (MIRO (Cha et al., 2022)), or output space (CAR-FT (Mao et al., 2024),CLIPood (Shu et al., 2023)).
+The general idea can be applied to weight space (L2SP (Xuhong et al., 2018)), feature space (MIRO (Cha et al., 2022)), or output space (e.g., CAR-FT (Mao et al., 2024), CLIPood (Shu et al., 2023)).
 Large-Scale Pretraining for DG: Recent DG literature (Cho et al., 2023; Cha et al., 2022; Addepalli et al., 2024; Mao et al., 2024; Arpit et al., 2022) leverages large-scale pretrained initializations stronger than ImageNet (Russakovsky et al., 2015), and CLIP (Radford et al., 2021) is the most common choice.
@@
 (2024) investigated how the presence of perceptually similar images in CLIP (Radford et al., 2021) pretraining affects performance, introducing the Similarity Hypothesis, which posits that nearest neighbor similarity is strongly correlated with zero-shot accuracy.
+Mayilvahanan et al.
+(2025) show that domain contamination in the pre-training has substantial impact on DG performance.
 Udandarao et al.
@@
 In contrast, we examine the domain generalization setting, where pre-trained models are fine-tuned on source domains and tested on held-out target domains.
+Yu et al.
+(2024) recommend using self-supervised pre-training to avoid data label leakage.
+In contrast we study DG model behavior in the more realistic setting of CLIP-pretraining.
 Our findings suggest that comparing target images to pre-trained images, as proposed by  Mayilvahanan et al.
@@
 3.1 Testing the Image Similarity Hypothesis
+Figure 2: Comparing the Predictive Power of the Alignment and Image Similarity Hypotheses for Domain Generalization (DG).
+a.) Image Similarity Hypothesis: The cosine similarity between a test image and its closest match from the pre-training set (Perceptual Similarity Score) shows relatively weak predictive power for accuracy, implying that visual resemblance alone is not fully indicative of downstream performance.
+b.) Alignment Hypothesis: In contrast, the cosine similarity between image and ground truth text-label embedding after pre-training (Alignment Score) is highly predictive of model accuracy after fine-tuning on five DG datasets, with Alignment Score distributions shown in the colored histograms.
+This suggests that image-text pairs well-aligned during pre-training result in better performance on target tasks.
 The Image Similarity Hypothesis (Mayilvahanan et al., 2024) posits that test performance improves when there are perceptually similar images in the pre-training dataset.
@@
 […]
-Figure 2: Comparing the Predictive Power of the Alignment and Image Similarity Hypotheses for Domain Generalization (DG).
-a.) Alignment Hypothesis: The cosine similarity between image and ground truth text-label embeddings before pre-training (Alignment Score) is highly predictive of model accuracy after fine-tuning on five DG datasets, with Alignment Score distributions shown in the colored histograms.
-This suggests that image-text pairs well-aligned during pre-training result in better performance on target tasks.
-b.) Image Similarity Hypothesis: In contrast, the cosine similarity between a test image and its closest match from the pre-training set (Perceptual Similarity Score) shows weaker predictive power for accuracy, implying that visual resemblance alone is less indicative of downstream performance.
 […]
+Figure 2 a.) shows the results of this analysis of the recent, high-performing VL2V-SD (Addepalli et al., 2024) on various DG datasets.
+While the Image Similarity Hypothesis is somewhat predictive of DG performance, its influence is not very strong.
+This suggests that perceptually similar pretraining data alone may not guarantee high DG performance; additional factors, such as how effectively the nearest neighbors were aligned with the target concept, may also be significant.
+Algorithm 1 Evaluating the Image Similarity Hypothesis
+1:Target domain samples $D_{\text{target}}$, trained DG model $M$, pre-trained image encoder $f_{I}$,
+2:for each sample $I\in D_{\text{target}}$ do
+3:     Retrieve nearest neighbor of $I$ in LAION-400M using $f_{I}$ features, assign to $I_{k}$
+4:     Compute PerceptualSimilarityScore($I,I_{k}$) using Equation Eq. 1,
+5:     Record correctness of $M(I)$
+6:end for
+7:Bin samples based on PerceptualSimilarityScore
+8:Compute DG accuracy within each bin
+9:return Accuracy for each bin
 3.2 Introducing the Alignment Hypothesis
@@
 More formally:
-$\text{AlignmentScore}(I,T)=\frac{\langle f_{I}(I),f_{T}(T)\rangle}{\|f_{I}(I)\|\cdot\|f_{T}(T)\|}$
+$\text{AlignmentScore}(I,T)=\frac{\langle f_{I}(I),f_{T}(T)\rangle}{\|f_{I}(I)\|\cdot\|f_{T}(T)\|}$ (Eq. 2)
 where $f_{I}(I)$ is the embedding of the image before finetuning on source, and $f_{T}(T)$ is the embedding of the text.
+Relative to directly using the contrastive loss value which also depends on negatives, PerceptualSimilarityScore has a scale which aligns across datasets (Appendix B.6).
 We verify the Alignment Hypothesis similarly to the Image Similarity Hypothesis, by binning samples using the AlignmentScore and computing accuracy for each bin using VL2V-SD.
 We provide the same analysis for many more DG methods in the Appendix Figure 11.
-In Figure 2 a.), we can see that the Alignment Hypothesis explains DG performance after source finetuning, significantly more strongly than for the Image Similarity Hypothesis in Figure 2 b.) This finding suggests that source fine-tuning in DG, which aims to achieve high performance across all target samples, only succeeds on those with high initial alignment.
+In Figure 2 b.), we can see that the Alignment Hypothesis explains DG performance after source finetuning, significantly more strongly than for the Image Similarity Hypothesis in Figure 2 a.) This finding suggests that source fine-tuning in DG, which aims to achieve high performance across all target samples, only succeeds on those with high initial alignment.
 4 Re-thinking Domain Generalization Benchmarking using the Alignment Hypothesis
@@
 We start by computing AlignmentScore for all samples in 5 DG datasets (Figure 3).
-Based on our observation that some samples are mislabeled, we perform dataset cleaning (4.1).
-Finally, we find an AlignmentScore threshold to split DG datasets into In-Pretraining (IP) and Out-of-Pretraining (OOP) evaluation subsets (Section 4.2), which we later use for evaluating DG methods (Section 5).
+Based on our observation that some samples are mislabeled, we perform dataset cleaning (Section 4.1).
+Then we find an AlignmentScore threshold to split DG datasets into well aligned In-Pretraining (IP) and poorly aligned Out-of-Pretraining (OOP) evaluation subsets (Section 4.2), which we later use for evaluating DG methods (Section 5).
 In order to connect the AlignmentScore with the DG method, we use the same backbone both for splitting the datasets into IP and OOP subsets and for training DG methods.
@@
 At very low scores, we find that a large fraction of labels are incorrect (red boxes in Figure 3).
-As a result, we divide the data into AlignmentScore intervals (e.g., 0.00-0.05, 0.05-0.10, and so on, up to 0.2) and randomly sample 100 instances from each interval for every dataset.
+Thus, we divide the data into AlignmentScore intervals (e.g., 0.00-0.05, 0.05-0.10, and so on, up to 0.2) and randomly sample 100 instances from each interval for every dataset.
 This allows us to systematically analyze the relationship between AlignmentScore and label accuracy across different score ranges.
@@
 Therefore, we discard all samples with AlignmentScore less than 0.15 in DomainBed-IP/OOP.
-As shown in Table 1, we observe that the percentage of discarded samples due to mislabeling varies across datasets, with VLCS and DomainNet having the highest rates at 12.41% and 7.64%, respectively.
-Furthermore, on the right side of Figure 3, we observe that at very high AlignmentScores (greater than 0.4), the images often contain text directly related to the label.
-Since our goal is to evaluate visual recognition rather than text recognition (OCR), and CLIP is known to have strong OCR abilities  (Fort, 2021), we exclude all samples above AlignmentScore of 0.4 from DomainBed-IP/OOP.
-As shown in Table 1, although only a small portion of data is removed due to OCR filtering (0.00-0.15% across datasets), this issue may become more significant in future studies.
-Table 1: Percentage of discarded samples due to mislabeling or label text in the image (OCR).
-Dataset Dropped - Mislabelled Dropped - OCR
-OfficeHome 0.92% 0.15%
-PACS 3.05% 0.00%
-TerraIncognita 0.22% 0.00%
-DomainNet 7.64% 0.03%
-VLCS 12.41% 0.00%
+As shown in Table 8 (in the Appendix), we observe that the percentage of discarded samples due to mislabeling varies across datasets, with VLCS and DomainNet having the highest rates at 12.41% and 7.64%, respectively.
+Furthermore, on the right side of Figure 3, we observe that at very high AlignmentScores (greater than 0.4), images often contain text directly related to the label.
+However, our goal is to evaluate visual recognition rather than text recognition (OCR), and CLIP is known to have strong OCR abilities (Fort, 2021), so we exclude all samples above AlignmentScore of 0.4 from DomainBed-IP/OOP.
+As shown in Table 8 (in the Appendix), only a small portion of data is removed due to OCR filtering (0.00-0.15% across datasets), but this issue may become more significant in future studies.
 4.2 Data Splitting
 After filtering, we focus on determining a threshold to split the dataset into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets.
-We select 0.21 as the threshold, based on the trends observed in Figure 2 (a.)), as this is the point where performance begins to improve significantly, indicating that existing methods become more effective.
+We select 0.21 as the threshold, based on the trends observed in Figure 2 b.), as this is the point where performance begins to improve significantly, indicating that existing methods become more effective.
 While this threshold represents a somewhat subjective choice informed by observed patterns, we provide AlignmentScores in the released data, allowing researchers the flexibility to experiment with their own thresholds.
-Figure 4 illustrates how this split impacts the size and composition of each dataset.
-Certain domains naturally fall into the IP subset.
-For example, in DomainNet, the clipart domain is predominantly categorized as IP, likely due to its frequent presence on the internet and therefore the web-scraped pre-training data.
-On the other hand, TerraIncognita-OOP has relatively more balanced domains, but exhibits substantial class shift between OP and OOP splits (Figure 5), meaning some classes are better aligned than others during pre-training.
-More dataset statistics are in the Appendix.
-Figure 4: DomainBed-(IP/OOP) Statistics: Breakdown of DomainBed-IP and DomainBed-OOP counts, by dataset and domain.
-Overall, DomainNet and VLCS have the largest fraction of samples falling into DomainBed-OOP.
-Figure 5: Class-distribution shift: TerraIncognita’s class distribution differs between DB-IP and DB-OOP, indicating that some classes were better aligned during pretraining.
+Figure 18 in the Appendix shows how this split impacts the size and composition of each dataset.
+For example, clipart in DomainNet is predominantly categorized as IP, likely due to its frequent presence on the internet and, therefore, web-scraped pre-training data.
+In comparison, TerraIncognita-OOP is more balanced domains, but exhibits substantial class shift between IP and OOP splits (Figure 17 in the Appendix), meaning some classes are better aligned than others during pre-training.
+We further investigate the distinctions between the IP and OOP subsets using VisDiff (Dunlap et al., 2024), an LLM-based system that identifies differences between sets of images.
+For each combination of dataset, domain, and class—except for DomainNet, where we subsample roughly 15% of the combinations due to computational constraints—we independently sample up to 30 images from both the IP and OOP subsets.
+VisDiff employs CLIP (Radford et al., 2021) to compute an AUC-ROC score for the natural language differences proposed by the LLM, and we retain only those differences with an AUC score of 0.7 or higher.
+Results are presented in Appendix Table LABEL:tab:visdiff.
+Several clear patterns emerge from our analysis.
+Contextual or environmental elements often overshadow the primary object indicated by the text label.
+For example, in VLCS’s SUN09 domain, the OOP subset for the car class frequently features images described as “historical architecture,” while in Office Home the real-domain bed images are deemed OOP if they include “child-themed decor.” These findings imply that object size may also play a role.
+In the car example, visible architecture implies that the car occupies only a small portion of the image, and in the bed example, the presence of child-themed decor indicates that the entire bedroom is visible rather than just the bed.
+To test this hypothesis, we computed the object size of the ground truth labeled objects using bounding boxes generated by the open-world object detector OWLv2 (Minderer et al., 2023).
+We found object size is correlated with the AlignmentScore (see Appendix Figure 10), with the sole exception being TerraIncognita.
+This indicates that even the presence of small wildlife can substantially increase the AlignmentScore, and that the background vegetation does not present a significantly conflicting signal that is sometimes present in other datasets.
+Overall, our findings suggest that pre-training which better represents scenes with multiple objects or concepts may improve benchmark performance.
 5 Experiments
-5.1 Training and Evaluation Protocol
-We aim to evaluate how well Domain Generalization (DG) methods perform across both pretraining-aligned(IP) and pretraining-misaligned(OOP) data.
-We adhere to the DomainBed evaluation methodology, where one domain is chosen as the target, and the remaining domains act as source domains.
+We evaluate Domain Generalization (DG) performance across both pretraining-aligned (IP) and pretraining-misaligned (OOP) data.
+We adhere to the DomainBed evaluation methodology, where one domain is chosen as the target, and the remaining act as source domains.
 To maintain a sufficient amount of training data, we train all DG methods on the original, unsplit datasets.
 We use hyper-parameter values recommended by the original implementation authors for each method.
-After training, models are evaluated separately on the IP and OOP subsets, as well as the original, unsplit test domain.
-This approach allows us to measure how well each method generalizes to both pretraining-aligned and pretraining-misaligned data.
+After training, models are evaluated separately on IP and OOP subsets, as well as the original, unsplit test domain.
+This allows us to measure how well each method generalizes to both pretraining-aligned and pretraining-misaligned data.
 We follow the literature’s standard practice of computing performance per target domain and averaging the results across all domains.
 This method ensures that any domain imbalances do not disproportionately influence the final performance metrics.
-Table 2: Benchmarking DG methods on DB-IP/OOP
+5.1 Algorithms
+Explicit Regularization towards Pretraining: Several recent DG methods leverage explicit regularization towards the initialization.
+These methods generally operate either in weight space (by regularizing or freezing model parameters) or in feature space (by aligning internal feature representations with those of the pretrained model).
+MIRO (Cha et al., 2022) minimizes the Mutual Information between DG model intermediate features and CLIP intermediate features.
+Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022) freezes all parameters except those in the Multiheaded-Attention Layers.
+VL2V-SD (Addepalli et al., 2024) self-distills a linear combination of CLIP vision and text outputs into a model.
+CLIPood (Shu et al., 2023) regularizes both weights and output features.
+Weights are averaged between the pre-trained CLIP model and the fine-tuned model, and outputs are regularized using a loss that incorporates information from the pre-trained text encoder.
+Linear Probe - Fine Tuning (LP-FT) (Kumar et al., 2021) freezes the backbone, trains a linear probe, and then performs full finetuning.
+An untrained linear probe can cause finetuning to update the frozen backbone needlessly, potentially unlearning discriminative features.
+This biases the model towards pre-trained weights with smaller gradient updates.
+Domain Invariance: A classic idea for Domain Generalization is Domain Invariance in the feature space, where the model learns only class-discriminative features shared among all training/source domains.
+CORAL (Sun & Saenko, 2016) matches the second moments of features across different domains and has been shown to be highly effective (Gulrajani & Lopez-Paz, 2020).
+Flat Optima: Studies by  Izmailov et al.
+(2018) and  Cha et al.
+(2021) have shown that flat minima generalize better than sharp minima, as they make loss values less sensitive to perturbations in the loss surface, resulting in smaller increases in loss during domain shifts.
+SWAD (Cha et al., 2021) Averages model parameters during training, determining the interval over which to average using validation loss over source domain.
+Model Parameter Averaging (MPA) (Arpit et al., 2022) Starts averaging model parameters after a number of burn-in steps to find flat minima.
+SAGM (Wang et al., 2023) is an optimizer that explicitly optimizes for flat minima.
+Baseline and Oracle: We also train a baseline and oracle model for lower-bound and upper-bound reference.
+The baseline model is an Empirical Risk Minimization (ERM) model that is finetuned on source domains and evaluated on target domains, and has been found to be effective for the DG task (Gulrajani & Lopez-Paz, 2020).
+The oracle model is trained on an 80% training split of all domains and evaluated on a 20% test split.
+The oracle model removes the OOD aspect of generalization and provides a reasonable upper bound for DG methods.
+Finally, we add a zero-shot LAION-400M (Schuhmann et al., 2021) pre-trained ViT-B/16 from OpenClip, which is the pre-trained model used in our analysis.
+5.2 Results
+Table 1(c) reports results for the poorly pre-training aligned DomainBed-OOP, higly pre-training aligned DomainBed-IP, and standard DomainBed datasets.
+Underlined results highlight the best performance of any single method (excluding method combinations), while the bold numbers show the highest performance overall, excluding the upper bound and zeros-hot baseline.
+Overall, we find that AlignmentScore is highly predictive of performance.
+More detailed observations about the results follow, while a comparison to PerceptualSimilarityScore can be found in Appendix B.5.
+DG methods perform well on DomainBed-IP: In most datasets within DomainBed-IP, domain generalization approaches achieve excellent performance.
+On two out of five datasets ( OfficeHome and PACS), the best DG method even outperforms the oracle.
+A notable exception is TerraIncognita, where CLIPood scores 30% below the oracle, highlighting this dataset as a challenging outlier.
+Interestingly, all three datasets where performance exceeds the oracle have an average IP AlignmentScore of 0.28, while the others have a lower average AlignmentScore of around 0.26.
+Therefore, the underperformance of TerraIncognita may be partially explained by its lower IP AlignmentScore, suggesting that alignment plays a significant role in DG performance, even within the IP case.
+Another interesting observation is that on the IP subset, the zero-shot model can achieve performances greater than the finetuned models (for PACS and VLCS).
+This means that, for IP-data, DG finetuning sometimes causes more catastrophic forgetting than learning of new features We show additional analysis of zero-shot models in Appendix B.9.
+DG Methods leave much to be desired on DomainBed-OOP, but are still stronger than ERM: In DomainBed-OOP, we observe that even the top-performing DG methods struggle with low-alignment data.
+For example, CLIPood achieves 57.1% accuracy, which is a significant drop compared to its performance on DomainBed-IP (84.7%) and DomainBed-All (78.1%).
+Despite this, DG methods still outperform Empirical Risk Minimization (ERM) by up to 10% on DomainBed-OOP.
+Thus, DG methods are better equipped to handle domain shifts than ERM, possibly due to weak transfer of knowlegdge from pre-trained features.
+Nevertheless, there is still substantial room for improvement on low-alignment samples.
+Table 1: Benchmarking DG methods on DB-IP/OOP.
+Domain generalization methods excel on high-alignment (IP) datasets—often even surpassing the oracle—while their performance noticeably drops on low-alignment (OOP) data, though still outperforming ERM.
 DomainBed-IP DomainNet OfficeHome PACS TI VLCS Average
+OpenCLIP ZS 74.9 88.9 98.5 36.8 95.9 79.0
 CORAL (Sun & Saenko, 2016) 63.3 76.1 84.3 42.9 86.5 70.6
@@
 Upper Bound (Target Finetune) 81.6 88.5 97.8 93.4 93.8 91.0
-(a) Samples with high AlignmentScore values, indicating good alignment between image and text features learned during pretraining.
+(a) Samples with high AlignmentScore values, indicating good pretraining alignment.
 DomainBed-OOP DomainNet OfficeHome PACS TI VLCS Average
+OpenCLIP ZS 26.3 48.1 81.4 4.5 80.2 48.1
 CORAL (Sun & Saenko, 2016) 22.3 42.6 74.1 16.0 74.0 45.8
@@
 DomainBed-All DomainNet OfficeHome PACS TI VLCS Average
+OpenCLIP ZS 59.5 85.4 97.0 33.2 82.4 71.5
 CORAL (Sun & Saenko, 2016) 50.6 73.2 83.2 39.6 78.5 65.0
@@
 (c) Performance of DG methods on unsplit DomainBed
-5.2 Algorithms
-Explicit Regularization towards Pretraining: Several recent DG methods leverage explicit regularization towards the initialization.
-These methods generally operate either in weight space (by regularizing or freezing model parameters) or in feature space (by aligning internal feature representations with those of the pretrained model).
-  • MIRO (Cha et al., 2022): Minimizes the Mutual Information between DG model intermediate features and CLIP intermediate features.
-  • Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022): Freezes all parameters except those in the Multiheaded-Attention Layers.
-  • VL2V-SD (Addepalli et al., 2024): Self-distills a linear combination of CLIP vision and text outputs into a model.
-  • CLIPood (Shu et al., 2023): Regularizes both weights and output features.
-Weights are averaged between the pre-trained CLIP model and the fine-tuned model, and outputs are regularized using a loss that incorporates class information from the pre-trained text encoder.
-  • Linear Probe - Fine Tuning (LP-FT) (Kumar et al., 2021): Freezes the backbone, trains a linear probe, and then performs full finetuning.
-An untrained linear probe can cause finetuning to update the frozen backbone needlessly, potentially unlearning discriminative features.
-This biases the model towards pre-trained weights with smaller gradient updates.
-Domain Invariance: A classic idea for Domain Generalization is Domain Invariance in the feature space, where the model learns only class-discriminative features shared among all training/source domains.
-CORAL (Sun & Saenko, 2016) matches the second moments of features across different domains and has been shown to be highly effective (Gulrajani & Lopez-Paz, 2020).
-Flat Optima: Studies by Izmailov et al.
-(2018) and Cha et al.
-(2021) have shown that flat minima generalize better than sharp minima, as they make loss values less sensitive to perturbations in the loss surface, resulting in smaller increases in loss during domain shifts.
-  • SWAD (Cha et al., 2021): Averages model parameters during training, determining the interval over which to average using validation loss over source domain.
-  • Model Parameter Averaging (MPA) (Arpit et al., 2022): Starts averaging model parameters after a number of burn-in steps to find flat minima.
-  • SAGM (Wang et al., 2023): An optimizer that explicitly optimizes for flat minima.
-Baseline and Oracle: We also train a baseline and oracle model for lower-bound and upper-bound reference.
-The baseline model is an Empirical Risk Minimization (ERM) model that is finetuned on source domains and evaluated on target domains, and has been found to be effective for the DG task (Gulrajani & Lopez-Paz, 2020).
-The oracle model is trained on an 80% training split of all domains and evaluated on a 20% test split.
-The oracle model removes the OOD aspect of generalization and provides a reasonable upper bound for DG methods.
-5.3 Results
-We show results in Table 2(c) for DomainBed-OOP, DomainBed-IP, and standard DomainBed datasets.
-Underlined results represent the best performance of any single method (excluding method combinations), while the bold numbers show the highest performance overall, excluding the upper bound.
-DG methods have nearly solved DomainBed-IP: In most datasets within DomainBed-IP, domain generalization approaches have nearly solved the problem of generalization.
-On three out of five datasets (DomainNet, OfficeHome, and PACS), the best DG method even outperforms the oracle!
-A notable exception is TerraIncognita, where CLIPood scores only 63.5%, far below the oracle’s 93.4%, highlighting this dataset as a challenging outlier.
-Interestingly, all three datasets where performance exceeds the oracle have an average IP AlignmentScore of 0.28, while the others have a lower average AlignmentScore of around 0.26.
-Therefore, the underperformance of TerraIncognita may be partially explained by its lower IP AlignmentScore, suggesting that alignment plays a significant role in DG performance, even within the IP case.
-DG Methods leave much to be desired on DomainBed-OOP, but are still stronger than ERM: In DomainBed-OOP, we observe that even the top-performing DG methods struggle with low-alignment data.
-For example, CLIPood achieves 57.1% accuracy, which is a significant drop compared to its performance on DomainBed-IP (84.7%) and DomainBed-All (78.1%).
-Despite this, DG methods still outperform Empirical Risk Minimization (ERM), which scores only 47.0% on DomainBed-OOP.
-Therefore DG methods are better equipped to handle domain shifts than ERM, possibly due to weak transfer of knowlegdge from pre-trained features.
-Nevertheless, there is still substantial room for improvement on low-alignment samples.
-State-of-the-Art methods do not consistently outperform older methods on OOP data: While CLIPood (Shu et al., 2023) clearly outperforms other methods on DomainBed-All with 78.1% accuracy, it performs comparably to older methods on DomainBed-OOP.
-For instance, MIRO + MPA (Cha et al., 2022; Arpit et al., 2022) achieves 57.2% on DomainBed-OOP, which is nearly the same as CLIPood’s 57.1%.
-This suggests that CLIPood’s primary advantage comes from well-aligned samples in DomainBed-IP, where it reaches 84.7% accuracy, half a percent better than the next-best method.
-Model Parameter Averaging (MPA) boosts performance on OOP data: MPA performs significantly better than ERM on DomainBed-OOP, achieving 53.0% compared to ERM’s 47.0%.
-This is half a percent better than MIRO on OOP data, despite being 2% worse than MIRO on IP data.
-When combined with MIRO, MPA delivers the best performance on DomainBed-OOP, scoring 57.2%, slightly surpassing CLIPood.
+SOTA methods do not consistently outperform older methods on OOP data: While CLIPood (Shu et al., 2023) clearly outperforms other methods on DomainBed-All, it performs comparably to older methods on DomainBed-OOP.
+For example, MIRO + MPA (Cha et al., 2022; Arpit et al., 2022) achieves 57.2% on DomainBed-OOP, which is nearly the same as CLIPood’s 57.1%.
+This suggests that CLIPood’s primary advantage comes from well-aligned samples in DomainBed-IP, where obtains 0.5% better performance than the next-best method.
+Model Parameter Averaging (MPA) boosts performance on OOP data: MPA obtains a significant 6% gain over ERM on DomainBed-OOP.
+MPA is 0.5% better than MIRO on OOP data despite MPA being 2% worse than MIRO on IP data.
+When combined with MIRO, MPA delivers the best performance on DomainBed-OOP, slightly surpassing CLIPood.
 This suggests that MPA can complement other regularization-based methods like MIRO.
-On DomainBed-IP, MIRO + MPA scores 84.1%, close to CLIPood’s 84.7%, demonstrating versatility across both high- and low-alignment data.
+On DomainBed-IP, MIRO + MPA is less than 1% away from CLIPood’s performance, demonstrating versatility across both high- and low-alignment data.
 Interestingly, SWAD underperforms MPA on DomainBed-OOP by 2%, despite being conceptually similar.
 We attribute this to selecting the averaging interval on the source data, which introduces overfitting to source domains.
+5.3 Discussion
+As an increasing number of works in the Domain Generalization sub-field leverage pre-trained CLIP models for Domain Generalization benchmarks, it is important to better characterize the impacts of pre-training on DG.
+We leave the reader with the following takeaways:
+Pre-training Alignment Predicts DG Performance: Our study demonstrates that pre-training alignment, measured as the cosine similarity between image and text embeddings, is a robust predictor of DG performance.
+This holds true even after source fine-tuning, highlighting that the quality of alignment achieved during pre-training has a significant impact on the generalization capability.
+Current DG Methods Exploit Pre-training Rather Than Learning New Features: Our findings reveal a large difference in the performance of DG methods between pretraining-aligned (IP) and pretraining-misaligned (OOP) data.
+While state-of-the-art methods achieve near-oracle performance on IP data, they struggle significantly on OOP data.
+This indicates that current methods primarily leverage on pre-trained features rather than learning new, generalizable features from source data.
+Consequently, their success is heavily tied to the quality of pre-training, rather than fine-tuning.
+Benchmarks Should Reflect Pre-training Reliance: The reliance on pre-trained alignment calls for a reevaluation of DG benchmarks.
+Existing benchmarks often aggregate results across all target data, masking the limitations of DG methods on low-alignment samples.
+To address this, we propose splitting evaluation datasets into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets.
+This provides a clearer picture of where DG methods succeed and where they fail.
+We hope that our proposed DomainBed-IP/OOP splits will guide the development of future methods that are better equipped to handle low-alignment data while maintaining performance on high-alignment samples.
 6 Conclusion
-In this paper, we systematically explore how Domain Generalization (DG) methods rely on pre-trained feature alignment from models like CLIP.
+We systematically explore how Domain Generalization (DG) methods rely on pre-trained feature alignment from models like CLIP.
 We hypothesize that the alignment between image and text embeddings during pre-training strongly predicts DG performance.
 Our experiments confirm this, showing that methods perform well on high-alignment samples (DomainBed-IP) but struggle on low-alignment data (DomainBed-OOP).
-While we focus on multi-source DG, we expect similar results in single-source finetuning due to the foundational role of pre-trained alignment.
 Notably, state-of-the-art methods like CLIPood perform near oracle-level on aligned data but see significant drops on misaligned samples.
@@
 While foundation models will continue to advance, there will always be specialized distributions where they fail.
+Therefore we take the stance that improved DG finetuning remains an important avenure of research.
 We hope these findings inspire further research into improving generalization on low-alignment data, pushing DG beyond reliance on pre-trained alignment.
+7 Acknowledgments
 DISTRIBUTION STATEMENT A.
@@
 A.1 Training and Evaluation Details
-Code is zipped in the supplementary.
-We use a slightly modified MIRO Cha et al.
-(2022) codebase for training and evaluation.
+We use a slightly modified MIRO (Cha et al., 2022) codebase for training and evaluation.
 We use leave-one-out evaluation, where a model is trained on all domains except the evaluation domain.
 We emphasize that we use DomainBed-IP and DomainBed-OOP as evaluation data only, models are trained on full datasets.
-For training, we use an OpenCLIP-ViT-B/16 Ilharco et al.
-(2021) trained on LAION-400M Schuhmann et al.
-(2021).
-We use default hyper-parameters as defined by  Cha et al.
-(2022).
+For training, we use an OpenCLIP-ViT-B/16 (Ilharco et al., 2021) trained on LAION-400M (Schuhmann et al., 2021).
+We use default hyper-parameters as defined by  (Cha et al., 2022).
 This includes a learning rate of 5e-5, weight decay of 0.0, a batch size of 32 per-domain, an Adam Optimizer, and no dropout for all methods.
@@
 For MIRO, we use use regularizer loss weight of 1.0.
-For CORAL, we use a CORAL regularizer weight of 1.0, following  Cha et al.
-(2021).
+For CORAL, we use a CORAL regularizer weight of 1.0, following  (Cha et al., 2021).
 For LP-FT, we train the linear probe for 600 steps before unlocking the full backbone.
 For Model Parameter Averaging, we burn in the training for 600 steps before averaging iterates.
-For VL2V-SD and CLIPood, we directly use the author’s implementation and hyper-parameters, except initializing with OpenCLIP Ilharco et al.
-(2021).
+For VL2V-SD and CLIPood, we directly use the author’s implementation and hyper-parameters, except initializing with OpenCLIP (Ilharco et al., 2021).
 A.2 Training Compute
@@
 We normalize for dataset size, so that no dataset dominates the count.
-In Figures 9 through 13, we plot these statistics independently per dataset, and find the trends consistent across datasets.
+In Figures 12 through 16, we plot these statistics independently per dataset, and find the trends consistent across datasets.
 B.2 Per-Dataset Benchmarking Results
-We expand Table 2(c) in the main paper into per-dataset results in Table 5 through 19.
+We expand Table 1(c) in the main paper into per-dataset results in Table 10 through 24.
 B.3 Similarity of Target to Pre-training
-To evaluate the Image Similarity Hypothesis, we retrieve the nearest neighbors from the Laion-400M dataset Schuhmann et al.
-(2021).
+To evaluate the Image Similarity Hypothesis, we retrieve the nearest neighbors from the Laion-400M dataset (Schuhmann et al., 2021).
 This raises the question of how similar the target domains are to the pre-training data and whether the source domains might be even more similar.
@@
 These findings reinforce that the usefulness of the DomainBed-IP/OOP split is not confined to a specific backbone.
+Table 2: Benchmark results of different backbones on OfficeHome and PACS datasets.
 Backbone OfficeHome - IP PACS - IP OfficeHome - OOP PACS - OOP
@@
 DinoV2 87.4 97.1 58.9 85.0
-Table 3: Benchmark results of different backbones on OfficeHome and PACS datasets.
+B.5 Splitting DomainBed using PerceptualSimilarityScore
+In Figure 2 a.), we show that the slope of the relationship of Top-1 Accuracy vs Perceptual Similarity Score is positive but shallow.
+This suggests that using PerceptualSimilarityScore as an alternative to AlignmentScore for splitting DomainBed would not be very effective.
+To further prove this point, we split at a PerceptualSimilarityScore value of 0.86 in Table 3.
+We can see the differences are not very large between OOP and IP, indicating that AlignmentScore is a better thresholding metric.
+Table 3: Splitting DomainBed by PerceptualSimilarityScore. The differences between IP and OOP with this split are much lower than with Alignment Score.
+Dataset PACS VLCS TerraIncognita OfficeHome DomainNet
+Perceptual IP 97.2 74.8 60.6 86.8 61.4
+Perceptual OOP 95.6 77.4 42.9 78.3 55.3
+B.6 Comparing AlignmentScore with zero-shot classification confidence score
+We also consider an alignment score which takes into account uncertainties, and compute a score using the confidence of the zero-shot classifier formed by the pre-trained CLIP model for each sample.
+Specifically, for a sample with ground truth class c, we calculate the softmax over the logits output by the zero-shot classifier, and use the resulting probability p(c) as the score.
+We refer to this as the Calibrated AlignmentScore and show the results in Figure 6.
+Although the score predicts generalization for both OfficeHome and DomainNet, the scores have different scales for different datasets.
+In contrast out AlignmentScore does align across datasets to a greater degree (Figure 7)
+Figure 6: Top-1 DG Accuracy vs calibrated alignment: We use the confidence of the zero-shot classifier formed by the pre-trained CLIP models an alignment measure.
+Although the score predicts generalization for both OfficeHome and DomainNet, the scores have different scales for different datasets.
+Figure 7: Top-1 DG Accuracy vs AlignmentScore: The AlignmentScore introduced in our work have scales which are comparable across different datasets.
+We explore the effect of averaging PerceptualSimilarityScore and AlignmentScore in Figure 8.
+We can see that there is not much of a compositional effect, so therefore we stick with AlignmentScore as our generalization predictor.
+Figure 8: Combining PerceptualSimilarity Score and AlignmentScore: We explore the effect of averaging PerceptualSimilarityScore and AlignmentScore.
+There is no visible additional signal from averaging.
+We therefore stick with AlignmentScore as our generalization predictor.
+B.7 Image Similarity Hypothesis for source data
+The main drawback of the Image Similarity hypothesis is that it does not consider how well the nearest perceptual neighbor is learned during pre-training.
+One reason for a sample being poorly learned during pre-training is that the pre-training caption is not very relevant to the DG task.
+Source data is unlikely to have this issue, since source and target domains share labels.
+Therefore it is interesting to ask how strongly correlated the PerceptualSimilarityScore is with DG accuracy when measured between source and target.
+Indeed , as seen in Figure 9, there is a strong correlation.
+However, simply using source-data to compute the PerceptualSimilarity results in an incomplete understanding of the relationship between target data and the training procedure, due to the lack of consideration of the pre-training.
+In fact, zero-shot models with NO learning from source are very performant (Table 1(c))
+Figure 9: PerceptualSimilarity to Source for OfficeHome. DG performance is correlated with similarity to source images.
+B.8 Training from scratch
+We focus on analyzing the generalization capabilities of CLIP-based DG methods, as these have been demonstrated to be the strongest in recent work.
+However,from-scratch experiments are valuable to measure how much finetuning on source data can learn, and so we also trained ResNet-50 models from scratch, using the MPA (Model Parameter Averaging) method and present the results in Table 4.
+We can see that performance is very low even on ALL DomainBed data (unsplit), further confirming that existing DG algorithms are poor at learning new knowledge from source data alone, and they rely on strong pre-training for good performance.
+Table 4: Training ResNet-50 from scratch. DG performance is very poor, showing DG methods rely on strong pre-training
+DomainNet OfficeHome PACS TI VLCS Avg
+ResNet-50 from Scratch 7.6 17.2 36.3 23.8 55.2 28.0
+Clip ViT-B/16 58.9 82.0 94.3 50.7 82.3 73.6
+B.9 Analysis of Zero-Shot Behavior
+To better understand the behavior of Domain Generalization (DG) methods, we compute confusion matrices for zero-shot (ZS) and In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets for the PACS dataset using CLIPood.
+The results are shown in Tables 5 and 6.
+Table 5: Confusion matrix for the IP subset of the PACS dataset.
+DG: Incorrect DG: Correct
+ZS: Incorrect 141 70
+ZS: Correct 179 8231
+Table 6: Confusion matrix for the OOP subset of the PACS dataset.
+DG: Incorrect DG: Correct
+ZS: Incorrect 98 62
+ZS: Correct 15 890
+The confusion matrices reveal key insights about the behavior of DG methods:
+  • For the IP subset, DG methods catastrophically forget a significant number of samples, flipping correct predictions made by zero-shot models to incorrect ones.
+This suggests that DG methods can even have a negative value for IP data.
+Nevertheless, performance remains strong.
+  • For the OOP subset, DG methods flip very few correct samples to incorrect ones.
+However, the state-of-the-art (SOTA) method, CLIPood, is still unable to correct the majority of incorrectly classified samples in this subset.
+Overall, in both IP and OOP cases, the results indicate that CLIPood and other DG methods rely heavily on pre-training.
+These findings underscore the limitations of current DG methods and highlight areas for improvement.
 Appendix C Additional DomainBed-(IP/OOP) Statistics and Analysis
 C.1 Constituent Datasets
-We split 5 datasets in DomainBed-(IP/OOP),( VLCS Fang et al.
-(2013), DomainNet Peng et al.
-(2019), OfficeHome Ganin et al.
-(2016), PACS Li et al.
-(2017), and TerraIncognita Beery et al.
-(2018)).
+We split 5 datasets in DomainBed-(IP/OOP),( VLCS (Fang et al., 2013), DomainNet (Peng et al., 2019), OfficeHome (Ganin et al., 2016), PACS (Li et al., 2017), and TerraIncognita (Beery et al., 2018)).
 Here we provide basic statistics of each.
@@
 PACS has 9991 sampes and 4 domains: arts, cartoon, photo, and sketch. There are 7 classes. The dataset shift is style.
-C.2 Mislabeling Rates
-In Table 4 we present the mislabelling rates at various AlingmentScore values.
+C.2 Mislabeling and OCR Rates
+In Table 7 we present the mislabelling rates at various AlignmentScore values. In Table 8, we present how much data is removed due to mislabelling or OCR filtering.
 Table 7: Mislabeling rates across different AlignmentScore ranges.
@@
 VLCS 33% 14% 9% 3%
+Table 8: Percentage of discarded samples due to mislabeling or label text in the image (OCR).
+Dataset Dropped - Mislabelled Dropped - OCR
+OfficeHome 0.92% 0.15%
+PACS 3.05% 0.00%
+TerraIncognita 0.22% 0.00%
+DomainNet 7.64% 0.03%
+VLCS 12.41% 0.00%
 C.3 Class Distribution of DomainBed-(IP/OOP):
-In Figures 14 through 28, we provide class distribution statistics of different datasets before splitting and in our IP and OOP splits.
+In Figures 19 through 33, we provide class distribution statistics of different datasets before splitting and in our IP and OOP splits.
 We find some interesting patterns.
@@
 Similarly, both PACS (DomainBed-OOP) and VLCS (DomainBed-OOP) are dominated by person.
+C.4 VisDiff differences between DB-IP and DB-OOP
+We further investigate the distinctions between the IP and OOP subsets using VisDiff (Dunlap et al., 2024), an LLM-based system that identifies differences between sets of images.
+For each combination of dataset, domain, and class—except for DomainNet, where we subsample roughly 15% of the combinations due to computational constraints—we independently sample up to 30 images from both the IP and OOP subsets.
+VisDiff employs CLIP to compute an AUC-ROC score for the natural language differences proposed by the LLM, and we retain only those differences with an AUC score of 0.7 or higher.
+The complete results are presented in Table LABEL:tab:visdiff.
+We find that contextual or environmental elements often overshadow the primary object indicated by the text label.
+For example, in the VLCS dataset’s SUN09 domain, the OOP subset for the car class frequently features images described as “historical architecture,” while in the Office Home dataset, real-domain bed images are deemed OOP if they include “child-themed decor.”
+C.5 Relationship between object size and AlignmentScore
+To test whether object size is correlated with AlignmentScore, we computed the object size of the ground truth labeled objects using bounding boxes generated by the open-world object detector OWLv2 (Minderer et al., 2023).
+Overall, object size is correlated with the AlignmentScore (see Figure 10), with the sole exception being the TerraIncognita dataset.
+This indicates that even the presence of small wildlife can substantially increase the AlignmentScore, and that the background vegetation does not present a significantly conflicting signal that is sometimes present in other datasets.
+Overall, this suggests that pre-training which better represents scenes with multiple objects or concepts may improve benchmark performance.
+(a) All datasets.
+(b) OfficeHome.
+(c) PACS.
+(d) DomainNet.
+(e) TerraIncognita.
+(f) VLCS.
+Figure 10: Overall relationship between ground truth object bounding box size and AlignmentScore across different DomainBed datasets.
+Overall the relationship is positive, indicating that object size plays a role in whether a data point is IP or OOP.
+The exception is TerraIncognita, indicating that camera trap backgrounds do not strongly conflict with the presence of the wildlife to be classified.
+See Section C.5.
+Table 9: VisDiff: Differences Identified for Image Sets.
+We use VisDiff (Dunlap et al., 2024), an LLM-based system for describing differences between images sets, to identify differences in between IP and OOP images.
+We do so independently for each dataset, domain, and class grouping.
+Distractors unrelated to the class seem to be an issue.
+For more details see Section C.4.
+Dataset_Domain_Class Description of OOP Subset
+VLCS_SUN09_car Historical architecture
+VLCS_SUN09_bird variety of landscapes
+VLCS_SUN09_person indoor water parks
+VLCS_Caltech101_dog text about adoption and rescue services
+VLCS_LabelMe_chair people walking on city streets
+VLCS_LabelMe_dog gatherings of people
+VLCS_Caltech101_bird paintings
+VLCS_Caltech101_person murals on buildings
+VLCS_VOC2007_dog inside homes with multiple people
+VLCS_Caltech101_car presence of branded vehicles
+VLCS_SUN09_dog cooking areas
+VLCS_VOC2007_car double decker buses
+VLCS_LabelMe_bird people near bodies of water
+VLCS_VOC2007_person equestrian activities
+VLCS_VOC2007_bird people engaging in activities
+terra_incognita_location_38_rabbit sunny daytime scenes
+terra_incognita_location_38_squirrel streams in wooded areas
+terra_incognita_location_46_opossum animal walking
+terra_incognita_location_100_dog muddy areas
+terra_incognita_location_46_raccoon cloud formations
+terra_incognita_location_43_dog streams in the forest
+terra_incognita_location_100_bird a dirt road in a garden
+terra_incognita_location_46_squirrel camera trap captures
+terra_incognita_location_38_bird night vision
+terra_incognita_location_46_empty middle of the night
+terra_incognita_location_100_rabbit absence of man-made objects
+terra_incognita_location_38_opossum screens with green backgrounds
+terra_incognita_location_46_dog crocodiles
+terra_incognita_location_46_rabbit shadows in nature
+terra_incognita_location_100_squirrel dirt roads
+terra_incognita_location_38_cat a black bear resting on a log
+terra_incognita_location_43_squirrel muddy areas
+terra_incognita_location_43_bobcat an image of a crocodile
+terra_incognita_location_43_bird deserts
+terra_incognita_location_38_coyote black bears
+terra_incognita_location_46_bobcat streams in a wooded area
+terra_incognita_location_46_coyote waterfalls
+terra_incognita_location_46_cat Wooded areas
+terra_incognita_location_43_coyote crocodiles
+terra_incognita_location_43_opossum boats in water
+terra_incognita_location_100_cat images of tapirs at night
+terra_incognita_location_100_raccoon natural landscape
+terra_incognita_location_43_cat animals resting
+terra_incognita_location_43_raccoon wildlife in water
+office_home_Art_Paper_Clip animals hanging
+office_home_Clipart_Webcam colorful backgrounds
+office_home_Product_Soda rows of multiple beverage containers
+office_home_Real_Speaker jbl portable bluetooth speakers
+office_home_Product_Drill yellow and black color scheme
+office_home_Real_Clipboards advertising signage
+office_home_Clipart_Fan close-up views
+office_home_Real_Bike road cycling
+office_home_Product_Lamp_Shade circular objects
+office_home_Clipart_Notebook textbooks with colorful covers
+office_home_Real_Sink brushed nickel finish
+office_home_Clipart_Keyboard singular computer components
+office_home_Art_Speaker toy-like objects
+office_home_Product_Radio bluetooth devices
+office_home_Real_Glasses space-related elements
+office_home_Clipart_Desk_Lamp asian clothing styles
+office_home_Art_Bottle earrings with a bottle and cork
+office_home_Clipart_Clipboards a promotional offer
+office_home_Real_Flowers butterflies
+office_home_Product_Fan air purifiers
+office_home_Clipart_Alarm_Clock cartoon characters
+office_home_Art_Backpack toys with wings
+office_home_Art_Computer illustrations
+office_home_Art_Shelf flowers in pots
+office_home_Product_Pen focus on packaging
+office_home_Product_Mouse colorful electronics on display
+office_home_Clipart_Postit_Notes notebooks on a table
+office_home_Product_Clipboards natural materials without additional items
+office_home_Product_Batteries set of six objects
+office_home_Clipart_Curtains feminine imagery
+office_home_Art_Pan eggs with faces drawn on them
+office_home_Clipart_Telephone 3D rendering
+office_home_Clipart_Monitor technology themes
+office_home_Product_Screwdriver close-up of a person engaged in an activity
+office_home_Clipart_Paper_Clip Red color prominence
+office_home_Product_Calendar octagon shapes
+office_home_Product_Couch close-up furniture shots
+office_home_Art_Helmet dragon’s head
+office_home_Clipart_File_Cabinet 3D rendering
+office_home_Clipart_Shelf commercial products
+office_home_Clipart_TV realistic images
+office_home_Real_Curtains a bathroom setting
+office_home_Art_Toys a small white dog
+office_home_Art_Glasses individual with headgear
+office_home_Product_Sneakers gel cushioning
+office_home_Real_Drill blue colored objects
+office_home_Clipart_Glasses audio equipment
+office_home_Clipart_Lamp_Shade black and white drawing
+office_home_Real_Bed child-themed decor
+office_home_Art_File_Cabinet people in a professional setting
+office_home_Product_Backpack focus on material texture
+office_home_Art_Printer office settings with unusual elements
+office_home_Real_Batteries emphasis on a specific brand
+office_home_Art_Soda cans of angry birds tropic soda
+office_home_Real_Trash_Can objects on wheels
+office_home_Product_Eraser packages explicitly labeled ‘pentel’
+office_home_Real_Shelf laundry room
+office_home_Clipart_Eraser book on a black background
+office_home_Clipart_ToothBrush distinctive backgrounds
+office_home_Product_Glasses sunglasses with red lenses
+office_home_Clipart_Calendar sanitation theme
+office_home_Clipart_Mop detailed human figures
+office_home_Art_Oven combination of domestic and clinical spaces
+office_home_Real_Lamp_Shade minimalist style
+office_home_Art_Fan digital artwork
+office_home_Art_Webcam futuristic design
+office_home_Art_Laptop fantasy or fictional themes
+office_home_Art_Bed enchantment
+office_home_Art_Curtains dark hair
+office_home_Real_Fan multiple ceiling fans
+office_home_Clipart_Pen cocktail shakers
+office_home_Art_Screwdriver blue and silver color palette
+office_home_Product_Keyboard gaming mouse included
+office_home_Real_Notebook colorful stationery
+office_home_Real_Couch contemporary living spaces
+office_home_Art_Table standing posture
+office_home_Art_Pen creative process
+office_home_Real_Mop clothes hanging
+office_home_Art_Scissors video game characters
+office_home_Clipart_Scissors abstract shapes
+office_home_Art_Keyboard concrete structures
+office_home_Real_Calendar printable charts with decorative backgrounds
+office_home_Art_Calculator decorative elements
+office_home_Art_Bike animal on a vehicle
+office_home_Product_Webcam a single computer monitor
+office_home_Clipart_Bottle household container
+office_home_Clipart_Computer a printer
+office_home_Real_Sneakers Black high top sneakers
+office_home_Clipart_Mouse monochromatic themes
+office_home_Art_Notebook objects with a vintage look
+office_home_Clipart_Bucket objects associated with liquids
+office_home_Product_Laptop abstract simplicity
+office_home_Art_Push_Pin digital art
+office_home_Real_Push_Pin objects placed against natural backgrounds
+office_home_Art_Ruler animal skulls
+office_home_Art_Couch cats
+office_home_Art_Desk_Lamp intricate costumes
+office_home_Art_Mouse unique artistic interpretations
+office_home_Art_Refrigerator a toy girl in an unusual setting
+office_home_Product_Knives survival or tactical items
+office_home_Product_Helmet an image of a protective gear
+office_home_Clipart_Helmet abstract concepts
+office_home_Product_Notebook calendar
+office_home_Product_Paper_Clip a caliper and a paper clip
+office_home_Product_Hammer hammer head detail
+office_home_Clipart_Bed detailed human emotions
+office_home_Clipart_Chair a red background
+office_home_Art_Flowers cakes with unique decorations
+office_home_Product_Toys White and brown plush toys
+office_home_Art_Fork red dress
+office_home_Real_Pan cooked fish
+office_home_Real_Bucket tabletop settings
+office_home_Art_Trash_Can drinks on a colorful background
+office_home_Real_Knives focus on object
+office_home_Real_Pencil pink color theme
+office_home_Product_File_Cabinet green objects on a white background
+office_home_Real_Hammer German craftsmanship
+office_home_Art_Sink illustrations
+office_home_Clipart_Screwdriver minimalistic design
+office_home_Art_Marker illustrations of animals
+office_home_Art_Telephone sport equipment
+office_home_Real_Folder blue wall
+office_home_Product_Folder focus on texture
+office_home_Clipart_Batteries objects on a plain, uncluttered background
+office_home_Product_Mug black handle
+office_home_Product_Flowers a single red rose
+office_home_Real_Soda mango juice with a straw
+office_home_Art_Lamp_Shade people with hats
+office_home_Clipart_Oven focus on a single object
+office_home_Art_Spoon fantasy elements
+office_home_Clipart_Table cartoons
+office_home_Clipart_Hammer food items
+office_home_Art_Radio yellow boomboxes
+office_home_Clipart_Sink human activity
+office_home_Clipart_Drill digital art style
+office_home_Art_Eraser illustrations or drawings
+office_home_Clipart_Exit_Sign clear and singular message
+office_home_Clipart_Pencil a spool of ribbon
+office_home_Art_Batteries graphic t-shirts
+office_home_Art_Mug coffee with food items inside
+office_home_Clipart_Bike signs with text
+office_home_Art_TV multiple tv screens
+office_home_Product_Chair inflatable furniture
+office_home_Art_Drill conceptual illustrations
+office_home_Art_Postit_Notes cookies
+office_home_Art_Chair fire or flames
+office_home_Art_Hammer women in fantasy attire
+office_home_Art_Folder men in suits
+office_home_Product_Computer MSI branded devices
+office_home_Product_Push_Pin geographic representations
+office_home_Clipart_Laptop a pair of devices
+office_home_Art_Sneakers octopuses on shoes
+office_home_Art_Candles whimsical objects
+office_home_Real_Chair ottomans
+office_home_Clipart_Pan stacked objects
+office_home_Art_Bucket man playing music
+PACS_cartoon_person images featuring ice cream
+PACS_art_painting_dog hunting scenes
+PACS_art_painting_giraffe deer in the artwork
+PACS_art_painting_guitar artwork displaying paint splatters
+PACS_art_painting_person artistic depictions featuring abstract or imaginative elements
+PACS_cartoon_horse giraffes
+PACS_cartoon_guitar hand-drawn illustrations
+PACS_photo_dog dogs in vehicles
+PACS_photo_horse non-living objects
+PACS_art_painting_horse bulls
+domain_net_clipart_shorts characters with speech bubbles
+domain_net_infograph_bottlecap technical illustrations
+domain_net_infograph_kangaroo different types of fish in infographics
+domain_net_clipart_bicycle scenes involving flight
+domain_net_sketch_sock monkeys as a subject
+domain_net_painting_stethoscope text with humor or messages
+domain_net_infograph_aircraft_carrier an organized visual representation of navy ships
+domain_net_clipart_The_Eiffel_Tower black background
+domain_net_sketch_submarine educational elements like drawing guides
+domain_net_clipart_parrot blue birds
+domain_net_painting_string_bean paintings on a striped background
+domain_net_real_house large resorts
+domain_net_real_stop_sign images near a river
+domain_net_clipart_mouse cute small animals
+domain_net_real_hedgehog objects in a grassy area
+domain_net_real_cat presence of dogs
+domain_net_painting_bucket watercolor paintings
+domain_net_sketch_motorbike steampunk design
+domain_net_sketch_wine_glass cocktails
+domain_net_painting_strawberry lace tablecloths
+domain_net_sketch_bus text elements in illustrations
+domain_net_infograph_flower floral arrangements
+domain_net_clipart_tiger Airplanes flying in the sky
+domain_net_clipart_lobster repeated phrases or words
+domain_net_sketch_baseball a basketball on a wooden floor
+domain_net_infograph_steak comparison charts
+domain_net_sketch_pillow teddy bear motifs
+domain_net_clipart_fish black and white illustrations
+domain_net_infograph_hockey_puck infographics focusing on player statistics
+domain_net_painting_bandage children in art
+domain_net_painting_harp circular tables
+domain_net_sketch_ant clouds
+domain_net_sketch_calculator squares and shapes
+domain_net_sketch_church specific locations
+domain_net_infograph_paintbrush informative visual aids
+domain_net_painting_sea_turtle stained glass
+domain_net_clipart_candle birthday-themed images
+domain_net_clipart_baseball_bat characters with beverages
+domain_net_painting_squirrel a combination of different artistic styles
+domain_net_infograph_trumpet mushrooms
+domain_net_real_cloud illustrations of devices
+domain_net_infograph_string_bean dairy farm flyer
+domain_net_clipart_scorpion designs for personal adornment
+domain_net_painting_mouth art creation process
+domain_net_painting_pickup_truck people creating art
+domain_net_painting_scorpion paintings with multiple subjects
+domain_net_clipart_asparagus chef’s hat
+domain_net_real_elephant frogs
+domain_net_painting_suitcase scenes involving rivers or water
+domain_net_real_tornado a video game setting
+domain_net_painting_pond a painting of people
+domain_net_painting_oven indoor car scenes
+domain_net_infograph_flip_flops modern info presentation template
+domain_net_clipart_vase casual outdoors
+domain_net_clipart_dolphin multiple animal species
+domain_net_real_lightning people indoors
+domain_net_infograph_banana recipes
+domain_net_painting_cooler incomplete words on objects
+domain_net_sketch_tractor depiction of people
+domain_net_clipart_power_outlet colorful graphic design
+domain_net_infograph_bandage infographics
+domain_net_real_string_bean recipes with bacon
+domain_net_infograph_sheep leather industry
+domain_net_real_carrot hummus
+domain_net_painting_lighthouse murals
+domain_net_clipart_skyscraper flyer designs
+domain_net_sketch_blueberry a set of hand drawn fruit
+domain_net_sketch_snorkel cartoon animals
+domain_net_infograph_vase flyers with promotional content
+domain_net_real_moon mythological or fantasy elements
+domain_net_sketch_hamburger fast food with ice cream
+domain_net_infograph_barn organic farming elements
+domain_net_clipart_rainbow umbrellas
+domain_net_clipart_dog images featuring human characters
+domain_net_infograph_elbow educational posters
+domain_net_clipart_face vector illustrations
+domain_net_sketch_hourglass vector illustrations
+domain_net_sketch_laptop people sitting at desks
+domain_net_real_fire_hydrant airplane in the sky
+domain_net_real_mouse a cable connection
+domain_net_real_hourglass jewelry
+domain_net_painting_whale paintings in a gallery setting
+domain_net_painting_envelope mixed media paintings
+domain_net_sketch_flip_flops illustrations of hats and sunglasses
+domain_net_infograph_hedgehog infographics with statistical data
+domain_net_painting_sun depictions of solitude
+domain_net_painting_knee military uniforms
+domain_net_clipart_butterfly birds and flowers
+domain_net_sketch_goatee illustrations of chefs
+domain_net_real_airplane dynamic outdoor sports
+domain_net_sketch_banana hand drawn illustrations
+domain_net_infograph_lollipop sweet-themed infographics
+domain_net_real_owl colorful designs
+domain_net_real_banana plate presentations
+domain_net_painting_tornado people in colorful dresses
+domain_net_real_lobster multiple dishes presented together
+domain_net_sketch_frog people drawing with different tools
+domain_net_real_mountain extreme sports
+domain_net_real_drums guitar
+domain_net_infograph_knee recovery processes
+domain_net_real_teapot coffee mugs
+domain_net_real_camouflage images with helicopters
+domain_net_infograph_parrot first aid themes
+domain_net_infograph_bathtub wheelchair-accessible features
+domain_net_painting_bracelet artistic portraits
+domain_net_painting_fireplace people sitting together
+domain_net_sketch_owl Rough outlines
+domain_net_sketch_crocodile lizards and frogs
+domain_net_clipart_kangaroo Christmas themes
+domain_net_clipart_cat cartoons with both cats and dogs
+domain_net_real_golf_club screenshots of video games
+domain_net_painting_laptop illustrated people
+domain_net_clipart_broom dog in a witch’s hat with a pumpkin
+domain_net_painting_bulldozer night scenes indoors
+domain_net_sketch_sailboat nautical-themed accessories
+domain_net_clipart_picture_frame Colorful paw prints
+domain_net_clipart_octagon colorful gumballs
+domain_net_clipart_ladder whimsical illustrations
+domain_net_clipart_scissors illustrations with human figures
+domain_net_clipart_tractor images of airplanes
+domain_net_real_chandelier rooms with large beds
+domain_net_sketch_guitar Everyday scenes involving musical themes
+domain_net_clipart_broccoli promotional style imagery
+domain_net_clipart_ice_cream refrigerators
+domain_net_painting_octagon birds
+domain_net_clipart_church ceremonial events
+domain_net_infograph_sleeping_bag child safety
+domain_net_painting_drill colorful outdoor scenes
+domain_net_painting_hot_tub green character in a bathroom scene
+domain_net_clipart_rhinoceros cartoon characters
+domain_net_sketch_roller_coaster people on a piece of paper
+domain_net_painting_bridge Venice
+domain_net_clipart_violin animals playing instruments
+domain_net_real_mosquito abstract or surreal elements
+domain_net_sketch_spider comic book grading
+domain_net_infograph_donut checkered patterns
+domain_net_painting_stairs art restoration
+domain_net_clipart_flamingo patterns and designs
+domain_net_real_knee people in stylish outfits
+domain_net_infograph_basketball detailed sports statistics
+domain_net_clipart_sleeping_bag colorful designs
+domain_net_painting_remote_control watercolor
+domain_net_painting_calculator a person multitasking
+domain_net_real_speedboat tourist beaches
+domain_net_clipart_bus psychedelic imagery
+domain_net_painting_speedboat images with a variety of painting mediums
+domain_net_painting_toilet decorative art
+domain_net_painting_cell_phone colorful designs
+domain_net_sketch_beard scenes depicting historical or fictional characters
+domain_net_clipart_key a man in a hat
+domain_net_clipart_spreadsheet vector illustrations with folders
+domain_net_infograph_mushroom identification charts
+domain_net_clipart_nose cartoon animal with a flower
+domain_net_painting_stop_sign oil paintings
+domain_net_sketch_baseball_bat vector illustrations on a black background
+domain_net_real_elbow colorful clothing
+domain_net_infograph_feather posters for events
+domain_net_clipart_dragon an airplane flying
 Figure 11: Plotting AlignmentScore vs Top-1 Accuracy for all benchmarked methods, for all datasets together.
@@
 Method Art Clipart Product Real Avg
-CORALSun & Saenko (2016) 33.0 29.4 50.0 58.0 42.6
-SAGMWang et al. (2023) 36.8 30.7 51.7 58.7 44.5
-ERM*Gulrajani & Lopez-Paz (2020) 36.0 28.1 51.3 56.0 42.9
-LP-FTKumar et al. (2021) 37.2 29.7 52.0 54.7 43.4
-SWADCha et al. (2021) 43.9 37.1 56.0 62.7 49.9
-MIRO Cha et al. (2022) 49.4 39.6 62.7 74.7 56.6
-VL2V-SD Addepalli et al. (2024) 56.4 39.3 65.3 65.3 56.6
-Attn. Tune Teterwak et al. (2023); Touvron et al. (2022) 46.0 36.1 59.7 64.0 51.4
-Model Parameter Averaging (MPA) Arpit et al. (2022) 46.0 36.4 55.7 66.0 51.0
-CLIPOOD (Shu et al., 2023) 60.3 44.4 73.3 78.0 63.9
+OpenClip ZS 49.4 24.6 59.0 59.3 48.1
+CORAL 33.0 29.4 50.0 58.0 42.6
+SAGM 36.8 30.7 51.7 58.7 44.5
+ERM 36.0 28.1 51.3 56.0 42.9
+LP-FT 37.2 29.7 52.0 54.7 43.4
+SWAD 43.9 37.1 56.0 62.7 49.9
+MIRO 49.4 39.6 62.7 74.7 56.6
+VL2V-SD 56.4 39.3 65.3 65.3 56.6
+Attn. Tune 46.0 36.1 59.7 64.0 51.4
+Model Parameter Averaging (MPA) 46.0 36.4 55.7 66.0 51.0
+CLIPOOD 60.3 44.4 73.3 78.0 63.9
 MIRO + SWAD 55.2 44.7 63.3 72.7 59.0
@@
 Method Art Clipart Product Real Avg
+OpenClip ZS 88.5 78.0 95.4 93.9 88.9
 CORAL 71.6 62.9 85.8 84.0 76.1
@@
 Method Art Clipart Product Real Avg
+OpenClip ZS 83.4 73.2 92.4 92.6 85.4
 CORAL 66.6 60.1 83.0 83.0 73.2
@@
 Model Parameter Averaging (MPA) 78.7 73.4 88.4 87.6 82.0
-CLIPOOD (Shu et al., 2023) 87.9 77.8 94.1 93.6 88.3
+CLIPOOD 87.9 77.8 94.1 93.6 88.3
 MIRO + SWAD 85.6 79.9 93.0 92.1 87.6
@@
 Method L100 L38 L43 L46 Avg
+OpenClip ZS 5.1 0.0 6.4 6.6 4.5
 CORAL 13.9 8.1 21.1 21.1 16.0
@@
 Model Parameter Averaging (MPA) 21.7 8.6 30.9 27.6 22.2
-CLIPOOD (Shu et al., 2023) 37.9 11.3 21.6 8.9 19.9
+CLIPOOD 37.9 11.3 21.6 8.9 19.9
 MIRO + SWAD 22.5 5.9 30.3 25.9 21.1
@@
 Method L100 L38 L43 L46 Avg
+OpenClip ZS 52.3 23.2 36.5 35.1 36.8
 CORAL 46.4 35.3 56.9 32.8 42.9
@@
 Model Parameter Averaging (MPA) 63.8 41.4 67.9 44.7 54.4
-CLIPOOD (Shu et al., 2023) 77.7 56.4 69.0 51.0 63.5
+CLIPOOD 77.7 56.4 69.0 51.0 63.5
 MIRO + SWAD 68.9 54.8 73.4 51.2 62.1
@@
 Method L100 L38 L43 L46 Avg
+OpenClip ZS 48.2 20.9 31.0 32.6 33.2
 CORAL 43.6 32.7 50.4 31.8 39.6
@@
 Model Parameter Averaging (MPA) 60.1 38.3 61.1 43.2 50.7
-CLIPOOD (Shu et al., 2023) 74.3 52.1 60.4 47.4 58.5
+CLIPOOD 74.3 52.1 60.4 47.4 58.5
 MIRO + SWAD 64.9 50.1 65.5 49.0 57.4
@@
 Method Clp Inf Pnt Qkdr Real Skt Avg
+OpenClip ZS 29.2 43.5 28.7 0.0 35.5 21.2 26.3
 CORAL 34.9 19.2 26.9 5.7 24.6 22.8 22.3
@@
 Model Parameter Averaging (MPA) 42.3 30.9 34.9 7.4 31.6 30.7 29.6
-CLIPOOD (Shu et al., 2023) 39.2 50.2 39.5 2.1 40.9 31.3 33.9
+CLIPOOD 39.2 50.2 39.5 2.1 40.9 31.3 33.9
 MIRO + SWAD 46.1 40.4 34.0 4.4 35.2 31.7 32.0
@@
 Method Clp Inf Pnt Qkdr Real Skt Avg
+OpenClip ZS 84.0 84.1 85.5 23.5 91.4 81.1 74.9
 CORAL 80.3 52.3 71.1 33.2 71.7 71.2 63.3
@@
 Model Parameter Averaging (MPA) 85.8 69.7 81.6 41.4 81.4 81.4 73.6
-CLIPOOD (Shu et al., 2023) 86.6 85.4 87.6 38.0 91.0 84.7 78.9
+CLIPOOD 86.6 85.4 87.6 38.0 91.0 84.7 78.9
 MIRO + SWAD 88.2 81.7 83.9 40.8 85.8 81.9 77.0
@@
 Method Clp Inf Pnt Qkdr Real Skt Avg
+OpenClip ZS 74.9 49.6 68.7 12.7 85.7 66.0 59.5
 CORAL 72.7 27.1 58.0 20.1 67.1 58.8 50.6
@@
 Model Parameter Averaging (MPA) 78.3 37.9 67.4 25.2 76.5 68.2 58.9
-CLIPOOD (Shu et al., 2023) 78.4 52.9 72.6 21.2 85.9 70.7 63.6
+CLIPOOD 78.4 52.9 72.6 21.2 85.9 70.7 63.6
 MIRO + SWAD 80.8 45.6 68.8 23.7 80.7 68.7 61.4
@@
 Method Caltech101 LabelMe SUN09 VOC2007 Avg
+OpenClip ZS 100.0 62.0 77.1 81.6 80.2
 CORAL 93.3 54.6 72.9 75.3 74.0
@@
 Model Parameter Averaging (MPA) 100.0 53.6 83.6 80.9 79.5
-CLIPOOD (Shu et al., 2023) 100.0 55.0 84.9 82.8 80.7
+CLIPOOD 100.0 55.0 84.9 82.8 80.7
 MIRO + SWAD 100.0 56.7 84.7 79.9 80.3
@@
 Method Caltech101 LabelMe SUN09 VOC2007 Avg
+OpenClip ZS 99.8 91.0 94.8 97.8 95.9
 CORAL 98.2 83.5 84.5 79.8 86.5
@@
 Model Parameter Averaging (MPA) 98.4 83.0 92.6 88.8 90.7
-CLIPOOD (Shu et al., 2023) 98.9 83.6 93.5 94.0 92.5
+CLIPOOD 98.9 83.6 93.5 94.0 92.5
 MIRO + SWAD 97.6 83.4 92.6 90.6 91.1
@@
 Method Caltech101 LabelMe SUN09 VOC2007 Avg
+OpenClip ZS 99.8 72.7 71.0 86.0 82.4
 CORAL 98.2 66.2 72.7 77.0 78.5
@@
 Model Parameter Averaging (MPA) 98.4 65.7 81.3 83.7 82.3
-CLIPOOD (Shu et al., 2023) 98.9 66.7 81.6 86.6 83.4
+CLIPOOD 98.9 66.7 81.6 86.6 83.4
 MIRO + SWAD 97.6 66.4 81.6 82.6 82.0
@@
 Method Art Cartoon Photo Sketch Avg
+OpenClip ZS 96.4 95.9 95.7 37.7 81.4
 CORAL 80.4 84.3 89.5 42.2 74.1
@@
 Model Parameter Averaging (MPA) 94.0 88.9 97.9 50.2 82.7
-CLIPOOD (Shu et al., 2023) 96.0 96.8 97.9 58.3 87.2
+CLIPOOD 96.0 96.8 97.9 58.3 87.2
 MIRO + SWAD 96.0 94.9 100.0 50.7 85.4
@@
 Method Art Cartoon Photo Sketch Avg
+OpenClip ZS 99.7 99.7 99.9 94.6 98.5
 CORAL 83.3 89.8 91.8 72.1 84.3
@@
 Model Parameter Averaging (MPA) 98.6 95.8 98.2 89.2 95.4
-CLIPOOD (Shu et al., 2023) 99.4 99.3 99.9 92.0 97.7
+CLIPOOD 99.4 99.3 99.9 92.0 97.7
 MIRO + SWAD 99.1 97.9 99.9 93.3 97.6
@@
 Method Art Cartoon Photo Sketch Avg
+OpenClip ZS 97.8 98.7 99.8 91.6 97.0
 CORAL 81.4 89.2 91.7 70.5 83.2
@@
 Model Parameter Averaging (MPA) 97.2 94.8 98.2 87.1 94.3
-CLIPOOD (Shu et al., 2023) 98.2 98.8 99.8 90.2 96.8
+CLIPOOD 98.2 98.8 99.8 90.2 96.8
 MIRO + SWAD 98.0 97.4 99.9 91.0 96.6
 MIRO + MPA 97.8 98.0 99.9 92.9 97.2
+Figure 17: DomainBed-(IP/OOP) Statistics: Breakdown of DomainBed-IP and DomainBed-OOP counts, by dataset and domain.
+Overall, DomainNet and VLCS have the largest fraction of samples falling into DomainBed-OOP.
+Figure 18: Class-distribution shift: TerraIncognita’s class distribution differs between DB-IP and DB-OOP, indicating that some classes were better aligned during pretraining.
 Figure 19: Class distribution of DomainNet-IP. Zoom in on pdf for best viewing.